### v2

# MnLargeSymbols’164 MnLargeSymbols’171

Abstract.
We show that the labeled Thompson groups and the twisted Brin–Thompson groups are boundedly acyclic. This allows us to prove several new embedding results for groups. First, every group of type $F_{n}$ embeds quasi-isometrically into a boundedly acyclic group of type $F_{n}$ that has no proper finite index subgroups. This improves a result of Bridson [9] and a theorem of Fournier-Facio–Löh–Moraschini [21, Theorem 2]. Second, every group of type $F_{n}$ embeds quasi-isometrically into a $5$-uniformly perfect group of type $F_{n}$. Third, using Belk–Zaremsky’s construction of twisted Brin–Thompson groups, we show that every finitely generated group embeds quasi-isometrically into a finitely generated boundedly acyclic simple group.
Key words and phrases:
Labeled Thompson groups, twisted Brin–Thompson groups, topological full groups, finiteness property, boundedly acyclic, uniformly perfect, embedding of groups.
2020 Mathematics Subject Classification:
57M07, 21J06
Introduction
Bounded cohomology of groups was first studied by Johnson and Trauber in the seventies in the text of Banach algebras [28]. The theory was then generalized to topological spaces in Gromov’s seminal work [24]. Since then, bounded cohomology has found applications in many different fields [13, 15, 25, 37], and has become an independent and active research field.
Though the bounded cohomology of a group can be defined rather easily (see for example [22, Chapter I]), it is hard to compute in general except in lower degrees. Surprisingly, Johnson proved that all amenable groups are boundedly acyclic [24, 28]. Recall that a group $G$ is boundedly acyclic if its bounded cohomology $\mathrm{H}_{b}^{i}(G,\mathbb{R})=0$ for any $i\geq 1$. This leads one to consider bounded acyclicity as a weak form of amenability for sometime. Besides amenable groups, Matsumoto and Morita proved in [31] that the groups of compactly supported homeomorphisms of Euclidean spaces are also boundedly acyclic. Similar techniques were developed to prove that mitotic groups [30] and binate groups [20] are boundedly acyclic. However, none of these non-amenable examples are finitely generated. Combining mitoses and suitable HNN extensions, Fournier-Facio, Löh, and Moraschini provided a way to functorially embed finitely generated groups into finitely generated boundedly acyclic groups [21, Theorem 2]. They also constructed the first finitely presented example [21, Theorem 4]. A breakthrough was then made by Monod [34], when he showed that the Lamplighter groups (i.e. any group of the form $G\wr\mathbb{Z}$) and the Thompson group $F$ are boundedly acyclic. Using new ideas in [34, 35], Andritsch was able to show that the Thompson group $V$ is also boundedly acyclic [2]. Note that the Thompson group $T$ however is not boundedly acyclic [20, Corollary 6.12]. More recently, Campagnolo, Fournier-Facio, Lodha and Moraschini provided further examples of boundedly acyclic groups in [16], including the finitely generated simple orderable groups constructed in [27, 29, 32]. In the current paper, we add a few new classes of boundedly acyclic groups to the list.
The first class of groups we consider is the labeled Thompson group $V(G)$ which has been independently defined before by Nekrashevych [40, Section 3.5] and Tanushevski in [45], see also [12]. The definition of $V(G)$ is similar to that of the Röver–Nekrashevych groups, except now the group $G$ acts trivially on the infinite rooted binary tree. In particular, $V(G)$ naturally surjects to $V$. Recall that a group is of type $F_{n}$ if it admits an aspherical CW-complex with fundamental group $G$ and finite $n$-skeleton, and a group is called $N$-uniformly perfect if every element can be written as the product of $N$ commutators. Uniformly perfect implies injectivity of the comparison map $H^{2}_{b}(-,\mathbb{R})\to H^{2}(-,\mathbb{R})$ [31, Corollary 2.11]. We summarize our construction in the following theorem.
Theorem 0.1.
There exists a functor $\mathbf{groups}\to\mathbf{groups}$ associating to each group $G$ the labeled Thompson group $V(G)$ with the following properties:
  1. (1)
    (1.13) There is an injective group homomorphism $\iota_{G}:G\to V(G)$ that is also a quasi-isometric embedding when $G$ is finitely generated;
  2. (2)
    (2.1) $V(G)$ has no proper finite index subgroups;
  3. (3)
    (2.4) $V(G)$ is $5$-uniformly perfect;
  4. (4)
    (3.14) $V(G)$ is of type $F_{n}$ if and only if $G$ is of type $F_{n}$;
  5. (5)
    (5.15) $V(G)$ is boundedly acyclic.
Remark 0.2.
One could also consider $V(G)$’s subgroup $F(G)$ whose finiteness property has been determined in [45] and [47, Section 6]. However $F(G)$ is not boundedly acyclic in general since it retracts to $G$, and it has many finite index subgroups since it surjects to the Thompson group $F$.
Remark 0.3.
More generally, given any injective wreath recursion $\phi$, we consider the corresponding $\phi$-labeled Thompson group $V_{\phi}(G)$ (see Section 1) and prove that it is always boundedly acyclic (Theorem 5.15). This includes for example the Röver–Nekrashevych groups corresponding to the infinite rooted binary tree. See also Corollary 6.4 for the general case.
We immediately have the following corollaries.
Corollary 0.4.
Every group of type $F_{n}$ embeds quasi-isometrically into a boundedly acyclic group of type $F_{n}$ that has no proper finite index subgroups.
Remark 0.5.
Without the boundedly acyclic property, this was established by Bridson [9]. Our construction also has the advantage that it is functorial.
Corollary 0.6.
Every group of type $F_{n}$ embeds quasi-isometrically into a $5$-uniformly perfect group of type $F_{n}$ that has no proper finite index subgroups.
The celebrated Higman’s embedding theorem asserts that every recursively presented group embeds into a finitely presented group. Combining this with Corollary 0.4, we have:
Corollary 0.7.
Every recursively presented group embeds into a finitely presented boundedly acyclic group that has no proper finite index subgroups.
Remark 0.8.
Without the no proper finite index subgroups property, this has been already established in [21, Corollary 5.2]. On the other hand, Baumslag, Dyer, Miller proved in [4, Theorem E] that every recursively presented group embeds into a finitely presented acyclic group.
Recall that Collins and Miller constructed a group of type $F$ (it is in fact of geometric dimension $2$) that has unsolvable word problem. Taking the labeled group $G$ to be their group, we have the following interesting application.
Corollary 0.9.
There exists a boundedly acyclic group of type $F_{\infty}$ that has unsolvable word problem and no proper finite index subgroups.
The second class of groups we consider is the twisted Brin–Thompson groups. These groups were first constructed by Belk and Zaremsky in [6] as a generalization of Brin’s high-dimensional Thompson groups [10]. In fact, given any group $G$ acting faithfully on a countable set $S$, they define the associated twisted Brin–Thompson group $\mathrm{SV}_{G}$ as a subgroup of the group of homeomorphisms of the Cantor space $\mathfrak{C}^{S}$. The key feature of twisted Brin–Thompson groups is that they are always simple [6, Theorem 3.4].
Theorem 0.10.
[Corollary 6.5] The twisted Brin–Thompson groups $\mathrm{SV}_{G}$ are all boundedly acyclic.
Remark 0.11.
By taking $G$ to be the Houghton group $\mathcal{H}_{n}$ and $S$ to be $n$ copies of natural numbers, we get an example of a boundedly acyclic simple group that is of type $F_{n-1}$ but not $F_{n}$ using [6, Corollary G].
More generally, we can show that topological full groups acting extremely proximally on the Cantor set are also boundedly acyclic, see 6.2. Combining 0.10 with [6, Theorem C], we have the following.
Corollary 0.12.
Every finitely generated group embeds into a finitely generated (indeed two-generated) boundedly acyclic simple group quasi-isometrically.
At this point, one might hope that all the $V$-like Thompson groups are boundedly acyclic. This however is not true, as Fournier-Facio, Lodha and Zaremsky showed in [19] that the braided $BV$ has infinite dimensional second bounded cohomology. It is also worth to pointing out that all the boundedly acyclic groups we have found are of infinite cohomological dimension. It is an interesting question whether there is a non-amenable boundedly acyclic group of type $F$.
Acknowledgements.
XW is currently a member of LMNS and supported by NSFC
No.12326601. He thanks Jonas Flechsig for discussions related to the finiteness properties of labeled Thompson groups. ZZ is partially supported by the Linghang Fundation. We thank Josiah Oh for proofreading part of the paper, Francesco Fournier-Facio for many useful comments on the paper, Volodymyr Nekrashevych for helpful communications on the history of $V_{\phi}(G)$. We also thank Matt Zaremsky for pointing out to us references about $V(G)$. In fact, after some communications with him, we realized that the labeled Thompson groups have already been defined and studied by Thompson in [46] under the name Splinter group (see also [17, Section 5.5]). In particular Corollary 2.2 has already proved been in [46, Theorem 2.3].
1. $\phi$-labeled Thompson groups
In this section, we introduce the basics of $\phi$-labeled Thompson groups, see also [40, Section 3.5] for more information. Note that the construction in [40, Section 3.5] works for general wreath recursions, but we will focus on injective wreath recursions for simplicity.
1.1. Wreath recursions
Let $\{0,1\}^{<\omega}$ be the monoid of finite words with letters in $\{0,1\}$ where the empty word is the unit and the multiplication is given by concatenation of words: $(x_{1}x_{2}\cdots x_{m})(y_{1}y_{2}\cdots y_{n})=x_{1}x_{2}\cdots x_{m}y_{1}y_{2}\cdots y_{n}$. So if $u\in\{0,1\}^{<\omega}$, then $u0,u1$ are the concatenations of $u$ by 00 and $1$. The length of a finite word $u$ is denoted by $l(u)$.
The infinite rooted binary tree $\mathcal{T}$ is the countably infinite rooted tree such that every vertex has degree $3$ except the root which has degree $2$. Note that $\{0,1\}^{<\omega}$ can be identified with the vertex set of $\mathcal{T}$, where the root corresponds to the empty word and edges are of the form $\{u,u0\}$ or $\{u,u1\}$ for any word $u$.
An automorphism of $\mathcal{T}$ is a bijection from $\{0,1\}^{<\omega}$ to $\{0,1\}^{<\omega}$ that fixes the root and preserves the edge relations, hence it also preserves the length of words. The group of automorphisms of $\mathcal{T}$ is denoted by $Aut(\mathcal{T})$.
We identify the Cantor set $\mathfrak{C}$ with $\{0,1\}^{\omega}$, the set of (right) infinite words with letters in $\{0,1\}$. Define $\iota:\{0,1\}^{\omega}\cup\{0,1\}^{<\omega}\to[0,1]$ by mapping $x_{1}x_{2}\cdots$ to $\Sigma_{i=1}^{\infty}\frac{x_{i}}{2^{i}}$ if the words have infinite length, and $x_{1}x_{2}\cdots x_{n}$ to $\Sigma_{i=1}^{n}\frac{x_{i}}{2^{i}}$ for finite words. Restricting to the subset of eventually 00 infinite words, $\iota$ is injective with image $\mathbb{Z}[\frac{1}{2}]\cap[0,1)$. We often identify the subset of eventually 00 infinite words in $\{0,1\}^{\omega}$ with $\mathbb{Z}[\frac{1}{2}]\cap[0,1)$ using $\iota$. Moreover, we have a lexicographical order on $\{0,1\}^{\omega}$ and $\{0,1\}^{<\omega}$ by declaring $0<1$, and $\iota$ is order preserving when restricted to eventually 00 infinite words. Note in general, $\iota$ is not injective even restricted to the set of finite words.
The concatenation of a finite word by an infinite word is defined by $(x_{1}x_{2}\cdots x_{m})(y_{1}y_{2}\cdots)=x_{1}x_{2}\cdots x_{m}y_{1}y_{2}\cdots$. Given a finite word $u$ and a (finite or infinite) word $v$, we say $u$ is a prefix of $v$ if $v=uw$ for some (finite or infinite) word $w$. The cone $u\{0,1\}^{<\omega}$ (resp. $u\{0,1\}^{\omega}$) of a finite word $u$ is defined as the subset of words in $\{0,1\}^{<\omega}$ (resp. $\{0,1\}^{\omega}$) which has $u$ as a prefix.
We will use $\mathrm{S}_{n}$ to denote the symmetric group of $[n]=\{1,\cdots,n\}$ where $\mathrm{S}_{n}$ acts on $[n]$ from the right. A wreath recursion of a group $G$ is a group homomorphism $\phi:G\rightarrow G\wr\mathrm{S}_{2}$, where $\mathrm{S}_{2}$ acts on $G\times G$ by permuting the coordinates. Let $\phi_{S}:G\rightarrow\mathrm{S}_{2}$ be the composition of $\phi$ with the natural projection $G\wr\mathrm{S}_{2}\rightarrow\mathrm{S}_{2}$. To make our life easier, we will focus on injective wreath recursions, i.e. when $\phi$ is injective.
Given a group $G$, there are some canonical injective wreath recursions one can consider:
Example 1.1.
  1. (1)
    The diagonal wreath recursion: $\Delta:g\rightarrow((g,g),id_{\mathrm{S}_{2}})$;
  2. (2)
    The right wreath recursion: $\phi_{r}:g\rightarrow{((1,g),id_{\mathrm{S}_{2}})}$;
  3. (3)
    $\phi_{\kappa}:g\rightarrow((g,g),\kappa(g))$ where $\kappa$ is a homomorphism from $G$ to $\mathrm{S}_{2}$.
Note that $Aut(\mathcal{T})$ has an canonical injective wreath recursion structure: the action of $Aut(\mathcal{T})$ on length-one words induces a homomorphism $\rho:Aut(\mathcal{T})\rightarrow\mathrm{S}_{2}$. Every element in $ker(\rho)$ can be regarded as an automorphism of the subtrees spanned by $0\{0,1\}^{<\omega}$ and $1\{0,1\}^{<\omega}$, hence $ker(\rho)$ is isomorphic to the direct product of two copies of $Aut(\mathcal{T})$ which $\mathrm{S}_{2}$ acts on by permuting the coordinates. This gives an isomorphism $\psi_{T}:Aut(\mathcal{T})\cong Aut(\mathcal{T})\wr\mathrm{S}_{2}.$
There is a natural (right) action of $G$ on the rooted binary infinite tree $\mathcal{T}$ with respect to $\phi$ The action of $G$ on $\mathcal{T}$ is defined by induction on the length of words:
  1. (1)
    $G$ preserves the empty word.
  2. (2)
    Suppose the action of words with length at most $n$ is defined. Then for $g\in G,v=xu$ where $l(u)=n,x\in\{0,1\}$, and $\phi(g)=((g_{0},g_{1}),\sigma_{g})\in G\wr\mathrm{S}_{2}$, we define $v\cdot g:=(x\cdot\sigma_{g})(u\cdot g_{x})$.
Therefore an injective wreath recursion $\phi$ gives a natural group homomorphism $\phi^{*}$ from $G$ to $Aut(\mathcal{T})$ such that the following diagram commutes:
[…]
Remark 1.2.
When $\phi^{*}$ is injective, i.e., the action is faithful, $G$ is called a self-similar group, see [39] for more information.
1.2. $G$-matrix and the $\phi$-labeled Thompson group.
Given an injective wreath recursion $\phi$ of a group $G$, we proceed to define the $\phi$-labeled Thompson group $V_{\phi}(G)$.
We call a finite subset $\Sigma\subset\{0,1\}^{<\omega}$ a partition set, if one of the following conditions holds (one easily checks these conditions are equivalent):
  1. (1)
    $u\{0,1\}^{<\omega}\cap v\{0,1\}^{<\omega}=\varnothing$ for any $u\neq v$ in $\Sigma$ and the set $\{0,1\}^{<\omega}-\bigcup_{u\in\Sigma}u\{0,1\}^{<\omega}$ is finite.
  2. (2)
    Any infinite word has an unique element in $\Sigma$ as its prefix.
  3. (3)
    There exists a number $n$ such that any finite word of length greater than $n$ has an unique element in $\Sigma$ as its prefix.
Every partition set $\Sigma$ gives a partition $\{u\{0,1\}^{\omega}:u\in\Sigma\}$ of $\{0,1\}^{\omega}$ which we can identify with the boundary $\partial\mathcal{T}$ of $\mathcal{T}$. Note that it also gives to a partition of the subset of eventually 00 infinite words (recall we identify it with $\mathbb{Z}[\frac{1}{2}]\cap[0,1)$). A partition of $\partial\mathcal{T}$ or $\mathbb{Z}[\frac{1}{2}]\cap[0,1)$ arising from a partition set is called a dyadic partition. The following lemma is immediate from the definition.
Lemma 1.3.
Let $u_{1},\cdots,u_{k}\in\{0,1\}^{<\omega}$ be distinct finite words. Then the cones $\{u_{i}\{0,1\}^{<\omega}\}_{i=1}^{k}$ are pairwise disjoint if and only if there is a partition set containing $\{u_{1},\cdots,u_{k}\}$ as a subset.
Now given $\sigma\in\mathrm{S}_{n}$ and partitions $\Sigma_{1}=\{u_{1},\cdots,u_{n}\},\Sigma_{2}=\{v_{1},\cdots,v_{n}\}$, we can define a homeomorphism of $\{0,1\}^{\omega}$:
$f^{\sigma}_{\Sigma_{1},\Sigma_{2}}:\{0,1\}^{\omega}\rightarrow\{0,1\}^{\omega},u_{i}w\mapsto v_{\sigma(i)}w.$
where $w$ is any infinite word. The subgroup of such homeomorphisms of $\{0,1\}^{\omega}$ is called the Thompson group $V$.
For a partition set $\Sigma$, if we replace one of its element $u$ by $\{u0,u1\}$ , we get a new partition set $\Sigma_{u}$, called the simple expansion of $\Sigma$ at $u$. We say a partition set $\Sigma_{1}$ is an expansion of another partition set $\Sigma_{2}$ if $\Sigma_{1}$ can be obtained from $\Sigma_{2}$ by finitely many simple expansions.
We proceed now to define the $\phi$-labeled Thompson group $V_{\phi}(G)$. A $G$-matrix is a block of the following form: $\scriptsize\begin{pmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{pmatrix}$ where $\{u_{1},\cdots,u_{n}\}$, $\{v_{1},\cdots,v_{n}\}$ are two partition sets and $g_{1},\cdots,g_{n}\in G$. $\{u_{1},\cdots,u_{n}\}$ is called the domain and $\{v_{1},\cdots,v_{n}\}$ is called the range of the matrix. We shall consider two $G$-matrices to be the same, if one is obtained from the other by permuting the columns. Note that we can always write a $G$-matrix in the form $\scriptsize\begin{pmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{pmatrix}$ where $u_{1}<\cdots u_{n}$ (here we are using the lexicographic order on $\{0,1\}^{<\omega}$). Note that the order of $\{v_{1},\cdots,v_{n}\}$ further gives a permutation $\sigma$, by mapping $i$ to the position of $v_{i}$ after we order them from small to big.
Definition 1.4 (Expansion rule).
Let $\phi:G\rightarrow G\wr\mathrm{S}_{2}$ be an injective wreath recursion. The simple expansion of the $G$-matrix $\scriptsize\begin{pmatrix}u_{1}&u_{2}&\cdots&u_{k}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{k}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{k}&\cdots&v_{n}\end{pmatrix}$ at $u_{k}$ is defined by:
$\scriptsize\begin{pmatrix}u_{1}&u_{2}&\cdots&u_{k}0&u_{k}1&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{k,0}&g_{k,1}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{k}(0)\sigma_{g_{k}}&v_{k}(1)\sigma_{g_{k}}&\cdots&v_{n}\end{pmatrix}$
where $((g_{k,0},g_{k,1}),\sigma_{g_{k}})=\phi(g_{k})$.
Given two $G$-matrices $\alpha,\beta$, we say $\beta$ is an expansion of $\alpha$ if it can be obtained by a sequence of simple expansions from $\alpha$. The inverse process of an expansion is called a reduction. We say two $G$-matrices are equivalent if they have a common expansion. The equivalence class of $\scriptsize\begin{pmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{pmatrix}$ is denoted by $\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}$.
Let $V_{\phi}(G)$ be the set of equivalence classes $G$-matrices with respect to $\phi$. We now define a multiplication on $V_{\phi}(G)$ to make it a group: For $a,b\in V_{\phi}(G)$, up to finite expansions, we can assume that the two $G$-matrices representing $a,b$ have the following forms: $a=\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}$ and $b=\scriptsize\begin{bmatrix}v_{1}&v_{2}&\cdots&v_{n}\\ h_{1}&h_{2}&\cdots&h_{n}\\ w_{1}&w_{2}&\cdots&w_{n}\end{bmatrix}$, now define $ab=\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}h_{1}&g_{2}h_{2}&\cdots&g_{n}h_{n}\\ w_{1}&w_{2}&\cdots&w_{n}\end{bmatrix}$.
Lemma 1.5.
The multiplication given above is well-defined, and makes $V_{\phi}(G)$ a group.
Proof.
We first show that the multiplication is well-defined. Note first that any two representatives of a $G$-matrix have a common expansion. Thus to prove the multiplication is well-defined, it suffices to prove that if two $G$-matrices $\alpha,\beta$ can perform multiplication already, and their expansions $\alpha^{\prime},\beta^{\prime}$ can also do, then $[\alpha\beta]=[\alpha^{\prime}\beta^{\prime}]$. Since expansions are compositions of finitely many simple expansions, we can further assume that $\alpha^{\prime},\beta^{\prime}$ are simple expansions of $\alpha$ and $\beta$.
Now given two $G$-matrices $a,b$, we can assume $\alpha=\scriptsize\begin{pmatrix}u_{1}&u_{2}&\cdots&u_{k}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{k}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{k}&\cdots&v_{n}\end{pmatrix}$ and $\beta=\scriptsize\begin{pmatrix}v_{1}&v_{2}&\cdots&v_{k}&\cdots&v_{n}\\ h_{1}&h_{2}&\cdots&h_{k}&\cdots&h_{n}\\ w_{1}&w_{2}&\cdots&w_{k}&\cdots&w_{n}\end{pmatrix}$. Let
$\alpha^{\prime}=\scriptsize\begin{pmatrix}u_{1}&u_{2}&\cdots&u_{k}0&u_{k}1&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{k,0}&g_{k,1}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{k}(0)\sigma_{g_{k}}&v_{k}(1)\sigma_{g_{k}}&\cdots&v_{n}\end{pmatrix},~{}\beta^{\prime}=\scriptsize\begin{pmatrix}v_{1}&v_{2}&\cdots&v_{k}0&v_{k}1&\cdots&v_{n}\\ h_{1}&h_{2}&\cdots&h_{k,0}&h_{k,1}&\cdots&h_{n}\\ w_{1}&w_{2}&\cdots&w_{k}(0)\sigma_{h_{k}}&w_{k}(1)\sigma_{h_{k}}&\cdots&w_{n}\end{pmatrix}$
be the simple expansions of $\alpha$ and $\beta$ at $u_{k}$ and $v_{k}$ where where $((g_{k,0},g_{k,1}),\sigma_{g_{k}})=\phi(g_{k})$ and $((h_{k,0},h_{k,1}),\sigma_{h_{k}})=\phi(h_{k})$. Then we have
$[\alpha^{\prime}][\beta^{\prime}]=\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{k}0&u_{k}1&\cdots&u_{n}\\ g_{1}h_{1}&g_{2}h_{2}&\cdots&g_{k,0}h_{k,(0)\sigma_{g_{k}}}&g_{k,1}h_{k,(1)\sigma_{g_{k}}}&\cdots&g_{n}\\ w_{1}&w_{2}&\cdots&w_{k}x_{0}&w_{k}x_{1}&\cdots&w_{n}\end{bmatrix}$
where $x_{i}=(i)\sigma_{g_{k}}\sigma_{h_{k}},i\in\{0,1\}$. Since $\phi(g_{k}h_{k})=((g_{k,0}h_{k,(0)\sigma_{g_{k}}},g_{k,1}h_{k,(1)\sigma_{g_{k}}}),\sigma_{g_{k}}\sigma_{h_{k}})$, we get $\alpha^{\prime}\beta^{\prime}$ is an expansion of $\alpha\beta$, hence $[\alpha^{\prime}][\beta^{\prime}]=[\alpha][\beta]$.
The associative law is obvious, and for any $a=\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}$ we have $\scriptsize\begin{bmatrix}v_{1}&v_{2}&\cdots&v_{n}\\ g^{-1}_{1}&g^{-1}_{2}&\cdots&g^{-1}_{n}\\ u_{1}&u_{2}&\cdots&u_{n}\end{bmatrix}$ as its inverse. This completes the proof that $V_{\phi}(G)$ is a group. $\square$
Definition 1.6.
The group $V_{\phi}(G)$ is called the $\phi$-labeled Thompson group. When $G$ is trivial, $V_{\phi}(G)$ is just the Thompson group $V$.
Remark 1.7.
Note that there is an induced action of $V_{\phi}(G)$ on the Cantor set $\{0,1\}^{\omega}$ via the almost automorphisms on $\mathcal{T}$: given $\alpha=\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}$, define $(u_{i}w)\alpha=v_{i}(w)\phi^{*}(g_{i})$ where $\phi^{\ast}$ is defined at the end of Subsection 1.1. This defines a right $V_{\phi}(G)$ action on $\mathfrak{C}$, we write $(w)\alpha$ as the image of $w$ under the action of $\alpha$.
Remark 1.8.
When $\phi^{\ast}$ in 1.1 is injective, $V_{\phi}(G)$ is known as the Röver–Nekrashevych group [38]. In [44] Skipper and Zaremsky prove that all Röver–Nekrashevych groups arise naturally from a cloning system [47]. In our setup, we have the flexibility that the action is not necessarily faithful. But one could also use the cloning system to define our $V_{\phi}(G)$. In fact $V_{\phi}(G)$ can be defined by the system of groups $G_{n}=G\wr\mathrm{S}_{n}$ with suitable cloning maps induced by $\phi$. In particular this gives another proof that $V_{\phi}(G)$ is a group. We leave the details to the reader.
1.3. Tree diagrams and G-matrices
We discuss in this subsection how one can represent elements in $\phi$-labeled Thompson group by $G$-labeled paired tree diagrams.
Recall that a finite rooted binary tree is a connected subtree of $\mathcal{T}$ containing the root, where every vertex has degree $1$ or $3$ except the root which has degree $2$ (or degree 0 if the root is also a leaf). A vertex with degree $1$ is called a leaf. A non-leaf vertex together with the two vertices directly below it is called a caret.
Observation 1.9.
The collection of partition sets of $\{0,1\}^{\omega}$ corresponds bijectively to the collection of finite rooted binary trees, the operation of simple expansions on partition sets corresponds to adding carets to the corresponding trees. See Figure 1.
00
$10$
$11$
00
$11$
$100$
$111$
expansion
expansion
Figure 1. An example of the correspondence of expansions.
To see this, for every partition set $\Sigma$, there is a unique finite rooted binary tree, denoted by $T_{\Sigma}$, such that $\Sigma$ is precisely the leaves ( i.e. vertices of degree one) in $T_{\Sigma}$. And for each finite rooted binary tree $T$, the set of leaves in $T$ is exactly a partition set, denoted by $\Sigma(T)$.
Given two finite rooted binary trees $T_{1},T_{2}$, we say $T_{1}$ is a simple expansion of $T_{2}$ if $T_{1}$ is obtained from the $T_{2}$ by adding a caret to one of the leaves; We say $T_{1}$ is an expansion of $T_{2}$ if $T_{1}$ is obtained from $T_{2}$ by applying finitely many simple expansions. Obviously $\Sigma_{1}$ is a (simple) expansion of $\Sigma_{2}$ if and only if $T_{\Sigma_{1}}$ is a (simple) expansion of $T_{\Sigma_{2}}$.
We can now represent elements of $V_{\phi}(G)$ by $G$-labeled paired tree diagrams. In fact, given a $G$-matrix $a=\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}$, there are two trees $T_{u}$ and $T_{v}$ representing the partition sets $\{u_{1},\cdots,u_{n}\}$ and $\{v_{1},\cdots,v_{n}\}$. Let $\sigma$ be the corresponding bijection of leaves that maps $u_{i}$ to $v_{i}$ for $1\leq i\leq n$, then we can write $a$ as $(T_{u},((g_{1},\cdots,g_{n}),\sigma)),T_{v})$. In Figure 2, we draw a $G$-labeled paired tree diagram with $T_{u}$ on the top and $T_{v}$ on the bottom where $\sigma$ is indicated by the arrow and $g_{i}$ is indicated by the labeling of the arrows.
00
$10$
$11$
$00$
$01$
$1$
Figure 2. An example of correspondence between $G$-matrix and tree diagram.
Similarly, one can define simple expansion and expansion for a $G$-labeled paired tree diagrams using the injective wreath recursion map $\phi$. And we shall denote the equivalence class of such diagrams under expansion by $[T_{u},((g_{1},\cdots,g_{n}),\sigma)),T_{v}]$. See Subsection 3.1 for more details.
1.4. The labeled Thompson group $V(G)$
Given a group $G$, recall that the diagonal wreath recursion $\Delta:G\rightarrow G\wr\mathrm{S}_{2}$, sends $g\mapsto((g,g),id_{\mathrm{S}_{2}})$. In this case, we shall simply write $V_{\Delta}(G)$ as $V(G)$, and call it the labeled Thompson group.
Given any two groups $G,H$, a group homomorphism $f:G\to H$ naturally induces a homomorphism $V(f):V(G)\rightarrow V(H)$, defined by sending $\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}$ to $\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ f(g_{1})&f(g_{2})&\cdots&f(g_{n})\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}$. Note that $V(-)$ defines a functor from the category of groups to itself.
Lemma 1.10.
The functor $V(-)$ maps surjections to surjections, injections to injections.
Proof.
The surjective part is clear from definition. Let $f:G\rightarrow H$ be an injective group homomorphism, and suppose that $x=\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}\in ker(V(f))$, then we have
$\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ f(g_{1})&f(g_{2})&\cdots&f(g_{n})\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}=1$
Since the identity element in $V(H)$ must be of the form $\scriptsize\begin{bmatrix}w_{1}&w_{2}&\cdots&w_{m}\\ 1_{H}&1_{H}&\cdots&1_{H}\\ w_{1}&w_{2}&\cdots&w_{m}\end{bmatrix}$ (up to expansion) for some partition set $\{w_{1},\cdots,w_{m}\}$, we conclude first that $f(g_{i})=1_{G}$, then $u_{i}=v_{i}$ for all $1\leq i\leq n$. But $f$ is injective, so $g_{i}=1_{G}$ for all $i$. Hence $x=1_{V(G)}$. $\square$
Another way to see $V(G)$ is to view it as a subgroup of a wreath product which we explain now. Recall that the Thompson group $V$ acts on $\mathfrak{C}$, let $G^{\mathfrak{C}}\rtimes V$ be the unrestricted wreath product of $G$ by $V$. Then we can define an embedding $i_{G}:V(G)\rightarrow G^{\mathfrak{C}}\rtimes V$ in the following way: for $a=\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ g_{1}&g_{2}&\cdots&g_{n}\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix}\mapsto(h_{a}:\mathfrak{C}\rightarrow G,\scriptsize\begin{bmatrix}u_{1}&u_{2}&\cdots&u_{n}\\ 1_{G}&1_{G}&\cdots&1_{G}\\ v_{1}&v_{2}&\cdots&v_{n}\end{bmatrix})$ where $h_{a}(x)=g_{i}$ for $x\in u_{i}\{0,1\}^{\omega}$.
Let $p_{G}:G\to\{1\}$ be the group homomorphism maps all elements in $G$ to $1$. We then have an induced map $V(p_{G}):V(G)\to V(\{1\})=V$. Denote kernel by $lim_{\mathcal{T}}G$, then we have a short exact sequence
$1\rightarrow lim_{\mathcal{T}}G\rightarrow V(G)\xrightarrow{V(p_{G})}V\rightarrow 1$
Note that one can also describe the kernel $\lim_{\mathcal{T}}G$ as the following: $G_{n}=\oplus_{2^{n}}G$ with injection maps $G_{n}\to G_{n+1}$ given by $(g_{1},g_{2}\cdots g_{2^{n}})\to(g_{1},g_{1},g_{2},g_{2},\cdots,g_{2^{n}},g_{2^{n}})$, then $\lim_{\mathcal{T}}G=\lim_{n\to\infty}G_{n}$. Furthermore we have a splitting for $V(p_{G})$, $V(i_{G}):V=V(\{1\})\rightarrow V(G)$ induced by the unique injection $i_{G}:\{1\}\to G$. This way we also get the splitting $V(G)=lim_{\mathcal{T}}G\rtimes V$.
1.5. The quasi-retract from $V(G)$ to $G$
We define a quasi-retract from $V(G)$ to $G$ which implies $G$ is of type $F_{n}$ if $V(G)$ is. Recall that a function $f:X\to Y$ between two metric spaces is called coarse Lipschitz if there exist constants $C,D>0$ so that
$d(f(x),f(x^{\prime}))\leq Cd(x,x^{\prime})+D\text{ for all }x,x^{\prime}\in X$
For example, any homomorphism between finitely generated groups is coarse Lipschitz with respect to the word metrics. A function $\rho:X\to Y$ is said to be a quasi-retraction if it is coarse Lipschitz and there exists a coarse Lipschitz function $\iota:Y\to X$ and a constant $E>0$ so that $d(\rho\circ\iota(y),y)\leq E$ for all $y\in Y$. If such a function $\rho$ exists, $Y$ is said to be a quasi-retract of $X$. In this case, it is not hard to show that $\iota$ must be a quasi-isometric embedding of $Y$ into $X$.
Theorem 1.11.
[1, Theorem 8] Let $G$ and $Q$ be finitely generated groups such that $Q$ is a quasi-retract of $G$ with respect to word metrics corresponding to some finite generating sets. Then if $G$ is of type $F_{n}$, so is $Q$.
To apply 1.11, we first define an embedding $\iota_{G}:G\to V(G)$ by mapping $g$ to $[T_{1},((g,1),\mathrm{id}),T_{1}]$ where $T_{1}$ is the caret. We also have a map $\rho_{G}:V(G)\to G$ by mapping an element $[T,((g_{1},\cdots,g_{n}),\sigma),S]$ in $V(G)$ to $g_{1}$. One checks that this is a well-defined map, but not a group homomorphism in general. We do have $\rho_{G}\circ\iota_{G}=\mathrm{id}_{G}$. Recall that $V$ embeds naturally in $V(G)$ as the subgroup of $V(G)$ with all elements have labeling $1$.
Lemma 1.12.
$V(G)$ is generated by $V$ and $\iota_{G}(G)$.
Proof.
Let $H$ be the subgroup generated by $V$ and $\iota_{G}(G)$. Given any element
$[T,((g_{1},\cdots,g_{n}),\sigma),S]$, we can write it as the product
$[T,((g_{1},\cdots,g_{n}),\mathrm{id}),T][T,((1,\cdots,1),\sigma),S].$
The right hand already lies in $V$, so it suffices to prove that elements of the form $[T,((g_{1},\cdots,g_{n}),\mathrm{id}),T]$ lies in $H$. Let $T_{r}$ be a finite rooted binary tree obtained by adding a finite rooted binary tree at the right leaf of $T_{1}$, then elements of the form $[T_{r},((g_{1},\cdots,g_{n}),\mathrm{id}),T_{r}]$ where $g_{i}=1$ for all but one $i$, are conjugate to elements of the form $[T_{r},((g,1,\cdots,1),\mathrm{id}),T_{r}]$ by elements in $V$. But $[T_{r},((g,1,\cdots,1),\mathrm{id})),T_{r}]=[T_{1},((g,1),\mathrm{id}),T_{1}]$, so elements of the form $[T_{r},((g_{1},\cdots,g_{n}),\mathrm{id})),T_{r}]$ where $g_{i}=1$ for all but one $i$ belong to $H$. Therefore elements of the form $[T_{r},((g_{1},\cdots,g_{n}),\mathrm{id})),T_{r}]$ also lie in $H$ since $H$ is closed under multiplication. In particular $[T_{1},((1,g),\mathrm{id}),T_{1}]\in H$ for any $g\in G$. Now the same argument before would imply all elements of the form $[T_{l},((g_{1},\cdots,g_{n}),\mathrm{id}),T_{l}]$ where $T_{l}$ is obtained from $T_{1}$ by adding a finite rooted binary tree at its left leave. Since any elements of the form $[T,((g_{1},\cdots,g_{n}),\mathrm{id}),T]$ can be written as a product of the form $[T_{r},((h_{1},\cdots,h_{n_{r}}),\mathrm{id}),T_{r}][T_{l},((t_{1},\cdots,t_{n_{l}}),\mathrm{id}),T_{l}]$, they must also lie in $H$. $\square$
Theorem 1.13.
The group $G$ is a quasi-retract of $V(G)$ when $G$ is finitely generated. In particular $\iota:G\to V(G)$ is an quasi-isometric embedding when $G$ is finitely generated.
Proof.
Fix a generating set $S_{G}$ for $G$, $S_{V}$ for $V$. By Lemma 1.12, $S_{V}$ and $\iota_{G}(S_{G})$ generates $V(G)$. We proceed to show that the map $\rho_{G}:V(G)\to G$ is coarse Lipschitz with respect to the word metric on $V(G)$ and $G$. Note that:
  1. (1)
    $\rho_{G}(x~{}\iota_{G}(s))\in\{\rho_{G}(g),\rho_{G}(g)s\}$ for all $s\in S_{G}$ and $x\in V(G)$, and
  2. (2)
    $\rho_{G}(yx)=\rho_{G}(x)$ for any $y\in V$ and $x\in V(G)$.
It follows that $\rho_{G}$ is nonexpanding and hence coarse Lipschitz. $\square$
Corollary 1.14.
If $V(G)$ has type $F_{n}$, so does $G$.
Proof.
When $n\geq 2$, this follows from 1.11 and 1.13. For $n=1$, suppose $G$ is not finitely generated, then $G$ can be written as a proper infinite increasing union $\bigcup_{i=1}^{\infty}G_{i}$. So $V(G)$ can also be written as a proper infinite increasing union $\bigcup_{i=1}^{\infty}V(G_{i})$ which implies $V(G)$ is not finitely generated. $\square$
Remark 1.15.
Our proof relies on the existence of the quasi-retraction map $\rho_{G}$ which does not exist for general injective wreath recursions $\phi$. In fact, if one takes $G$ to be the Grigorchuk group which is only finitely generated, the corresponding Röver–Nekrashevych group is of type $F_{\infty}$ [5].
2. Normal subgroups of $V(G)$ and its uniform perfectness
We study the normal subgroup structure and uniform perfectness of $V(G)$ in this section. Note that the normal subgroup structure of $F(G)$ has already been studied in [45]. Recall that we view $V$ as the subgroup of $V(G)$ consisting of all elements with trivial labeling.
Given $g\in G,u\in\{0,1\}^{<\omega}$, we define an element of $V(G)$ by $\Lambda_{u}(g)=\scriptsize\begin{bmatrix}u&u_{2}&\cdots&u_{n}\\ g&1&\cdots&1\\ u&u_{2}&\cdots&u_{n}\end{bmatrix}$ where $\{u,u_{2},\cdots,u_{n}\}$ is a partition set. The definition of $\Lambda_{u}(g)$ does not depend on the choice of $u_{2},\cdots,u_{n}$, thus is well-defined.
Theorem 2.1.
If $N$ is a normal subgroup of $V(G)$ such that $N\cap V$ is non-trivial, then $N=V(G)$. In particular, $V(G)$ has no proper finite index subgroups.
Proof.
Since $V$ is simple, $N\cap V$ is nontrivial and normal in $V$, we have $N\cap V=V$ hence $V\leq N$.
Let $\Lambda_{u}(g)=\scriptsize\begin{bmatrix}u&u_{2}&\cdots&u_{n}\\ g&1&\cdots&1\\ u&u_{2}&\cdots&u_{n}\end{bmatrix}$ and $q=\scriptsize\begin{bmatrix}u&u_{2}&\cdots&u_{n}\\ 1&1&\cdots&1\\ v&v_{2}&\cdots&v_{n}\end{bmatrix}$ for some $v\in\{0,1\}^{<\omega}$, then $q^{-1}\Lambda_{u}(g)q=\scriptsize\begin{bmatrix}v&v_{2}&\cdots&v_{n}\\ g&1&\cdots&1\\ v&v_{2}&\cdots&v_{n}\end{bmatrix}=\Lambda_{v}(g)$, so $\Lambda_{u}(g)^{-1}\Lambda_{v}(g)=(\Lambda_{u}(g)^{-1}q^{-1}\Lambda_{u}(g))q$ is in $N$ since $q\in V\leq N$ and $N$ is normal.
[…]
Since any element in $V(G)$ of the form $a=\scriptsize\begin{bmatrix}u^{\prime}_{1}&u^{\prime}_{2}&\cdots&u^{\prime}_{n}\\ g^{\prime}_{1}&g^{\prime}_{2}&\cdots&g^{\prime}_{n}\\ v^{\prime}_{1}&v^{\prime}_{2}&\cdots&v^{\prime}_{n}\end{bmatrix}$ can be written as $a=\Lambda_{u_{1}}(g^{\prime}_{1})\Lambda_{u_{2}}(g^{\prime}_{2})\cdots\Lambda_{u_{n}}(g^{\prime}_{n})p(a)$ where $p(a)=\scriptsize\begin{bmatrix}u^{\prime}_{1}&u^{\prime}_{2}&\cdots&u^{\prime}_{n}\\ 1&1&\cdots&1\\ v^{\prime}_{1}&v^{\prime}_{2}&\cdots&v^{\prime}_{n}\end{bmatrix}\in V$, we conclude that $V(G)$ is generated by elements of the forms $\Lambda_{r}(g)$ and $V$, which are all contained in $N$, hence $N=V(G)$. $\square$
Corollary 2.2.
$V(G)$ is perfect.
Proof.
The commutator subgroup $V(G)^{\prime}$ of $V(G)$ is a normal subgroup containing $(V)^{\prime}$, the latter is equal to $V$ since $V$ is simple and non-abelian. Hence $V(G)^{\prime}$ must equal to $V(G)$ by 2.1. $\square$
Recall that we have a projection map $V({p_{G}}):V(G)\to V$ induced by the map $p_{G}:G\to 1$. Let us abbreviate $V(p_{G})$ by $pr_{G}$. We improve 2.1 to the following.
Theorem 2.3.
Any proper normal subgroup of $V(G)$ is a subgroup of $lim_{\mathcal{T}}G$.
Proof.
Let $N$ be a normal subgroup of $V(G)$, then $pr_{G}(N)\trianglelefteq V$, hence $pr_{G}(N)=V$ or $1$.
Assume that $pr_{G}(N)=V$. Then $N$ has an element $g=\begin{bmatrix}u&\cdots\\ h&\cdots\\ v&\cdots\end{bmatrix}$ where $\{u,v\}$ is an anti-chain, i.e., $u\{0,1\}^{<\omega}\cap v\{0,1\}^{<\omega}=\varnothing$ and $u\{0,1\}^{<\omega}\cup v\{0,1\}^{<\omega}$ is a proper subset in $\{0,1\}^{<\omega}$. For any anti-chain $\{r,s\}$, let $q=\begin{bmatrix}u&v&\cdots\\ h&1_{G}&\cdots\\ r&s&\cdots\end{bmatrix}$, and $p=q^{-1}gq$, then $p$ has the form $p=\begin{bmatrix}r&\cdots\\ 1_{G}&\cdots\\ s&\cdots\end{bmatrix}$. Now for any $f\in G$,
$p^{-1}\Lambda_{r}(f)p=\begin{bmatrix}s&s_{2}&\cdots&s_{n}\\ 1_{G}&g^{-1}_{2}&\cdots&g^{-1}_{n}\\ r&r_{2}&\cdots&r_{n}\end{bmatrix}\begin{bmatrix}r&r_{2}&\cdots&r_{n}\\ f&1_{G}&\cdots&1_{G}\\ r&r_{2}&\cdots&r_{n}\end{bmatrix}\begin{bmatrix}r&r_{2}&\cdots&r_{n}\\ 1_{G}&g_{2}&\cdots&g_{n}\\ s&s_{2}&\cdots&s_{n}\end{bmatrix}=\Lambda_{s}(f)$
Hence $[p,\Lambda_{r}(f)]=\Lambda^{-1}_{s}(f)\Lambda_{r}(f)\in N$ for any anti-chain $\{r,s\}$. Let $\pi:G\rightarrow G/N$ be the quotient map, then since $\{r,s0\}$$\{r,s1\}$, and $\{s0,s1\}$ are anti-chains, we have
$\pi(\Lambda_{r}(f))=\pi(\Lambda_{s}(f))=\pi(\Lambda_{s0}(f))\pi(\Lambda_{s1}(f))=\pi(\Lambda_{r}(f))^{2}$
Hence $\pi(\Lambda_{r}(f))=1$ and $\Lambda_{r}(f)\in N$. Since $lim_{\mathcal{T}}G$ is generated by elements of the form $\Lambda_{r}(f)$, we have $lim_{\mathcal{T}}G\leq N$. Hence $N=V(G)$ since $pr_{G}(N)=V$.
So if $N$ is proper normal subgroup, $pr_{G}(N)=\{1\}$ which means that $N\leq lim_{\mathcal{T}}G$. $\square$
We further improve Corollary 2.2 to the following.
Theorem 2.4.
$V(G)$ is $5$-uniformly perfect.
Proof.
Let us switch to $G$-labeled paired tree diagram to represent elements in $V(G)$. For any element $[T_{-},((g_{1},\cdots,g_{n}),\sigma),T_{+}]\in V(G)$, we have $[T_{-},((g_{1},\cdots,g_{n}),\sigma),T_{+}]$
$=[T_{-},((1_{G},g_{2},\cdots,g_{n}),\mathrm{id}),T_{-}][T_{-},((g_{1},1_{G}\cdots,1_{G}),\mathrm{id}),T_{-}][T_{-},((1_{G},\cdots,1_{G}),\sigma),T_{+}].$
Notice that $[T_{-},((1_{G},\cdots,1_{G}),\sigma),T_{+}]$ is an element in Thompson group $V$ and $V$ is $3$-uniformly perfect [23, Corollary 6.6]. So $[T_{-},((1_{G},\cdots,1_{G}),\sigma),T_{+}]$ is a product of at most $3$ commutators in $V(G)$ since $V$ is a subgroup of $V(G)$.
Let $s_{12}=(12)\in\mathrm{S}_{n}$ and $\bar{s_{12}}=((1_{G},\cdots,1_{G}),s_{12})\in G\wr\mathrm{S}_{n}$, then we have an element $[T_{-},\bar{s_{12}},T_{-}]\in V(G)$. Note that
$[T_{-},((g_{1},1_{G}\cdots,1_{G}),\mathrm{id}),T_{-}]^{[T_{-},\bar{s_{12}},T_{-}]}=[T_{-},((1_{G},g_{1},1_{G},\cdots,1_{G}),\mathrm{id}),T_{-}].$
Thus to prove $[T_{-},((g_{1},\cdots,g_{n}),\sigma),T_{+}]$ can be written as products of $5$ commutators, it suffices to prove elements of the form $[T,((1_{G},g_{2},\cdots,g_{n}),\mathrm{id}),T]$ are commutators.
Let us prove now $v=[T,((1_{G},g_{2},\cdots,g_{n}),\mathrm{id}),T]$ is a commutator. Let $T^{\prime}$ be the tree obtained from $T$ by adding $(n-1)$ new carets such that each new caret always attaches to the left most leaf, then $v=[T^{\prime},((1_{G},\cdots,1_{G},g_{2},\cdots,g_{n}),\mathrm{id}),T^{\prime}]$. Let $T^{\prime\prime}$ be the tree obtained from $T$ by adding $(n-1)$ new carets on each leaf except the left most one, we have $v$ also equals to $[T^{\prime\prime},((1_{G},g_{2},g_{2},\cdots,g_{n},g_{n}),\mathrm{id}),T^{\prime\prime}]$. Now let $\alpha\in\mathrm{S}_{2n-1}$ with $\bar{\alpha}=((1_{G},\cdots,1_{G}),\alpha)\in G\wr\mathrm{S}_{2n-1}$ such that $\bar{\alpha}((1_{G},\cdots,1_{G},g_{2},\cdots,g_{n}),\mathrm{id})\bar{\alpha}^{-1}=((1_{G},g_{2},1_{G},g_{3},\cdots,1_{G},g_{n},1_{G}),\mathrm{id})$. Then we have $vav^{-1}a^{-1}=[T^{\prime\prime},((1_{G},1_{G},g_{2},1_{G},g_{3},\cdots,1_{G},g_{n}),\mathrm{id}),T^{\prime\prime}]$ where $a=[T^{\prime\prime},\bar{\alpha},T^{\prime}]$. Similarly pick $\beta\in\mathrm{S}_{2n-1}$ with $\bar{\beta}=((1_{G},\cdots,1_{G}),\beta)\in G~{}\wr~{}\mathrm{S}_{2n-1}$ such that
$\bar{\beta}((1_{G},1_{G},g_{2},1_{G},g_{3},\cdots,1_{G},g_{n}),\mathrm{id})\bar{\beta}^{-1}=((1_{G},\cdots,1_{G},g_{2},\cdots,g_{n}),\mathrm{id}).$
Then $b(vav^{-1}a^{-1})b^{-1}=[T^{\prime},((1_{G},\cdots,1_{G},g_{2},\cdots,g_{n}),\mathrm{id}),T^{\prime}]=v$, where $b=[T^{\prime},\bar{\beta},T^{\prime\prime}]$. Thus we have $v$ is a commutator and we finish the proof. $\square$
3. Finiteness properties of labeled Thompson groups
In this section we determine the finiteness properties of $V(G)$. The proof follows the standard approach to determine finiteness properties of Thompson-like groups using paired forest diagrams, see for example [14, 18, 42, 43, 44].
3.1. Labeled Thompson groups by paired tree diagrams
Recall in Subsection 1.3, we have explained how to define a $\phi$-labeled Thompson group using $G$-labeled paired-tree diagram. We expand this now to $G$-labeled paired-forest diagram. For convenience, we shall only discuss the case when the wreath recursion is the diagonal wreath homomorphism.
Recall that a finite rooted binary tree is a finite tree such that every vertex has degree 3, except the leaves, which have degree 1, and the root, which has degree 2 (or degree 0 if the root is also a leaf). We will drop the word finite whenever it is clear from the context. And we will draw such trees with the root at the top and the nodes descending from it, down to the leaves. A non-leaf vertex together with the two vertices directly below it is called a caret. If the leaves of a caret in $T$ are leaves of $T$ , we will call the caret elementary. A forest is a finite linearly ordered union of rooted binary trees. If each tree in the forest is either a single root or a caret, we call the forest elementary.
Definition 3.1.
A $G$-labeled paired-forest diagram is a triple $(F_{-},(\vec{g},\sigma),F_{+})$ consisting of forests $F_{-},F_{+}$ with the same number of leaves, say $n$, and $(\vec{g},\sigma)\in G\wr\mathrm{S}_{n}$. When $F_{-},F_{+}$ both have only one component, it is the $G$-labeled paired tree diagram.
We define now expansion and reduction of a $G$-labeled paired forest diagram. Given a $G$-labeled paired forest diagram $\left(F_{-},(\vec{g},\sigma),F_{+}\right)$, and some $1\leq i\leq n$, one replaces $F_{\pm}$ with forests $F_{\pm}^{\prime}$ obtained from $F_{\pm}$ by adding a caret to the two leaves labeled $i$ in $F_{\pm}$. Then replace $(\vec{g},\sigma)$ with $(\vec{g}^{\prime},\sigma^{\prime})\in G\wr S_{n+1}$, by “doubling” the $i^{\text{th }}$ coordinate in $\vec{g}$ and $\sigma^{\prime}$ is the induced bijiection which stays the same on the old leaves and maps the new left (resp. right) leave of $F_{-}^{\prime}$ to the new left (resp. right) leave of $F_{+}^{\prime}$. The triple $\left(F_{-}^{\prime},(\vec{g}^{\prime},\sigma^{\prime}),F_{+}^{\prime}\right)$ is called the expansion of $\left(F_{-},(\vec{g},\sigma),F_{+}\right)$. Conversely, $\left(F_{-},(\vec{g},\sigma),F_{+}\right)$ is a reduction of $\left(F_{-}^{\prime},(\vec{g}^{\prime},\sigma^{\prime}),F_{+}^{\prime}\right)$. Two paired forest diagrams are equivalent if one is obtained from the other by a sequence of reductions or expansions. The equivalence class of $(F_{-},(\vec{g},\sigma),F_{+})$ is denoted by $[F_{-},(\vec{g},\sigma),F_{+}]$. See Figure 3 for an example.
Figure 3. An example of equivalent tree diagrams in the labeled Thompson group $V(G)$
Figure 4. An example of the multiplication of two elements in the labeled group $V(G)$
We call a $G$-labeled paired forest diagram reduced if no reduction can be performed. One checks that every $G$-labeled paired forest diagram is equivalent to a unique reduced $G$-labeled paired forest diagram. We can define a groupoid structure on the equivalence classes of $G$-labeled forest diagrams just as all the Thompson-like groups. In fact, given $[F_{-},(\vec{g},\sigma),F_{+}],[F_{-}^{\prime},(\vec{g},\sigma^{\prime}),F_{+}^{\prime}]$ such that $F_{+}$ and $F_{-}^{\prime}$ have the same number of components. When $F_{+}=F_{-}^{\prime}$, we declare
$[F_{-},(\vec{g},\sigma),F_{+}]\ast[F_{-}^{\prime},(\vec{g}^{\prime},\sigma^{\prime}),F_{+}^{\prime}]=[F_{-},(\vec{g},\sigma)(\vec{g}^{\prime},\sigma^{\prime}),F_{+}^{\prime}]$
where the multiplication of $(\vec{g},\sigma)$ and $(\vec{g}^{\prime},\sigma^{\prime})$ uses the group structure on $G\wr\mathrm{S}_{n}$. In general, $F_{+}$ is not necessarily equal to $F_{-}^{\prime}$, but since they have the same number of components, we could always make them the same after finitely many expansions. Thus we could choose different representatives of $[F_{-},(\vec{g},\sigma),F_{+}]$ and $[F_{-}^{\prime},(\vec{g}^{\prime},\sigma^{\prime}),F_{+}^{\prime}]$ using expansions, and define the multiplication. See Figure 4. We call the equivalences classes of $G$-labeled paired forest diagrams equipped with such multiplication structure the $G$-labeled Thompson groupoid $\mathscr{F}(G)$.
Define further $\mathscr{F}_{m,n}(G)$ to be the full subgroupoid of $\mathscr{F}(G)$ consisting of $G$-labeled paired tree diagrams $(F_{-},(\vec{g},\sigma),F_{+})$ such that $F_{-}$ has $m$ components and $F_{+}$ has $n$ components. Note that $\mathscr{F}_{n,n}(G)$ is a group and $\mathscr{F}_{1,1}(G)$ is the labeled Thompson group $V(G)$.
3.2. Brown’s criterion and discrete Morse theory
To determine the finiteness property of $V(G)$, we will apply Brown’s criterion [11]. But we shall use a version that is combined with discrete Morse theory. Recall given a piecewise Euclidean cell complex $X$, a map $h$ from the set of vertices of $X$ to the integers is called a height function, if each cell has a unique vertex maximizing $h$. For $t\in\mathbb{Z}$, define $X^{\leq t}$ to be the full subcomplex of $X$ spanned by vertices $x$ satisfying $h(x)\leq t$. Similarly, define $X^{<t}$ and $X^{=t}$. The descending star $\operatorname{{\operatorname{St}}{\downarrow}}(x)$ of a vertex $x$ is defined to be the open star of $x$ in $X^{\leq h(x)}$. The descending link $\operatorname{{\operatorname{Lk}}{\downarrow}}(x)$ of $x$ is given by the set of “local directions” starting at $x$ and pointing into $\operatorname{{\operatorname{St}}{\downarrow}}(x)$. More details can be found in [7].
Theorem 3.2.
Let $\Gamma$ be a group acting cellularly on a contractible cubical complex $X$. Let $h:X\rightarrow\mathbb{R}$ be a $\Gamma$-equivariant height function. Given $m\geq 1$, then $\Gamma$ is of type $\mathrm{F}_{m}$ provided that the following conditions are satisfied:
  1. (1)
    For each $q\in\mathbb{R}$ the sublevel set $X^{h\leq q}$ is $\Gamma$-cocompact.
  2. (2)
    Each cell stabilizer is of type $\mathrm{F}_{m}$.
  3. (3)
    For each $n\leq m-1$ there exists $t\in\mathbb{R}$ such that for all $x\in X^{(0)}$ with $h(x)\geq t$ we have that the descending link $\operatorname{{\operatorname{Lk}}{\downarrow}}(x)$ is $n$-connected.
3.3. The $G$-labeled Stein–Farley complex
We build the $G$-labeled Stein-Farley complex that $V(G)$ acts on now. Note that the group $G\wr\mathrm{S}_{n}$ embeds in the groupoid $\mathscr{F}(G)$ by sending $(\vec{g},\sigma)$ to $[1_{n},(\vec{g},\sigma),1_{n}]$, where $1_{n}$ denotes the elementary forest consisting of only $n$ roots. With this embedding, we have a right action of $G\wr\mathrm{S}_{n}$ on $\mathscr{F}_{m,n}$ by right multiplication. Denote the orbit of an element $\alpha\in\mathscr{F}_{m,n}$ under this action by $[\alpha]$. Abuse of notation, we may sometimes denote the orbit of an equivalence class $[F_{-},(\vec{g},\sigma),F_{+}]$ also by $[F_{-},(\vec{g},\sigma),F_{+}]$. Note that if $\alpha\in\mathscr{F}_{m,n}$ and $\beta_{1},\beta_{2}\in\mathscr{F}_{n,l}$ with $[\alpha\ast\beta_{1}]=[\alpha\ast\beta_{2}]$, then $[\beta_{1}]=[\beta_{2}]$. We will refer to this as left cancellation.
Now let $\mathcal{P}(G)=\bigcup_{n=1}^{\infty}\mathscr{F}_{1,n}(G)/(G\wr\mathrm{S}_{n})$. We have a well-defined height function $h:\mathcal{P}(G)\rightarrow\mathbb{N}$ by sending elements in $\mathscr{F}_{1,n}(G)/(G\wr\mathrm{S}_{n})$ to $n$.
There is also a poset structure on $\mathcal{P}(G)$ which we explain now. Given $x,y\in\mathcal{P}(G)$ with $x=[T,(\vec{g},\sigma),F]$, say that $x\leq y$ if there exists a forest $F^{\prime}$ with $m$ leaves such that $y$ lies in the same orbit as $[T,(\vec{g},\sigma),F]\ast[F^{\prime},(\vec{1}_{G},\mathrm{id}),1_{m}]$ where $\vec{1}_{G}$ is the identity element in $G^{m}$. When this happens, $y$ is called a splitting of $x$. It is easy to check that this is a partial ordering on $\mathcal{P}(G)$. If the forest $F^{\prime}$ happens to be elementary, we denote the relation by $x\preceq y$ and call $y$ an elementary splitting of $x$. If $x\preceq y$ but $x\neq y$, then we write $x\prec y$. Note that $\preceq$ and $\prec$ are not transitive, though it is true that if $x\preceq z$ and $x\leq y\leq z$, then $x\preceq y$ and $y\preceq z$.
Note that $V(G)$ acts on $\bigcup_{n=1}^{\infty}\mathscr{F}_{1,n}(G)$ by left multiplication and this action commutes with right action of $G\wr\mathrm{S}_{n}$ and (elementary) splitting, hence $V(G)$ left acts on $\mathcal{P}(G)$ and preserves the relations $\leq$ and $\preceq$. Moreover the height function $h$ is $V(G)$-invariant.
Let $|\mathcal{P}(G)|$ be the geometric realization of the ordered space $(\mathcal{P}(G),\leq)$. Define $\mathcal{P}(G)^{\leq n}:=\{x\in\mathcal{P}(G):h(x)\leq n\}$, similarly $\mathcal{P}(G)^{=n}:=\{x\in\mathcal{P}(G):h(x)=n\}$ and $\mathcal{P}(G)^{\geq n}:=\{x\in\mathcal{P}(G):h(x)\geq n\}$. Then $|\mathcal{P}(G)^{\leq n}|$ is the full subcomplex of $|\mathcal{P}(G)|$ spanned by all the vertices $x$ with $h(x)\leq n$.
Definition 3.3 ($G$-labeled Stein–Farley complex).
Let $X_{G}$ be the subcomplex of $|\mathcal{P}(G)|$ consisting of all simplicies $[x_{0}<\cdots<x_{n}]$ satisfying $x_{0}\preceq x_{n}$.
Similarly, define $X_{G}^{\leq n}:=|\mathcal{P}(G)^{\leq n}|\cap X_{G}$. The complexes $X_{G}$, $|\mathcal{P}(G)^{\leq n}|,X_{G}^{\leq n}$ all admit a $V(G)$-action. The exactly same proof of [14, Proposition 2.1] shows that any two elements in $\in\mathcal{P}(G)$ have a least upper bound and if they have a common lower bound, they have a greatest lower bound. In particular, we have the following:
Lemma 3.4.
$(\mathcal{P}(G),\leq)$ is directed, i.e., every two elements in $\mathcal{P}(G)$ has a common upper bound with respect to the order $\leq$. Thus $|\mathcal{P}(G)|$ is contractible.
In $\mathcal{P}(G)$, for $x\leq y$ define the closed interval $[x,y]:=\{z\mid x\leq z\leq y\}$, and $|[x,y]|$ to be its geometric realization . Similarly, define the open and half-open intervals $(x,y),(x,y]$ and $[x,y)$. Note that if $x\preceq y$, then the closed interval $[x,y]$ is a Boolean lattice, and so the simplices in its geometric realization fit together into a cube. The top of the cube is $y$ and the bottom is $x$.
Now we want to show $X_{G}$ admits a cubical structure preserved by $V(G)$, where the cells are of the form $[x,y]$ for $x\preceq y$, and the function $t$ on the vertices of $X_{G}$ extends to a Morse function. Also, the intersection of cubes is either empty or is itself a cube; in fact if $[x,y]\cap[z,w]\neq\emptyset$, then $y$ and $w$ have a lower bound, and we get that $[x,y]\cap[z,w]=[\sup(x,z),\inf(y,w)]$ where $\sup(x,z)$ is the least upper bound of $x,z$ and $\inf(y,w)$ is the greatest lower bound of $y,w$. Therefore we can regard $X_{G}$ as a cube complex on which $V(G)$ acts cubically.
Proposition 3.5.
For each $n\geq 1$, $X_{G}^{\leq n}$ is $V(G)$-cocompact.
Proof.
Note first that for each $k\geq 1$, $V(G)$ acts transitively on $\mathcal{P}(G)^{=n}$. In fact, given two elements $[T_{1},a_{1},F_{1}]$ and $[T_{2},a_{2},F_{2}]$ in $\mathcal{P}(G)^{=n}$ we can first assume $F_{1}$ and $F_{2}$ are the same using the equivalence relation. Then $[T_{1},a_{1},F_{1}]=[T_{1},a_{1}a_{2}^{-1},T_{2}]\ast[T_{2},a_{2},F_{1}]$. Thus there exists for each $1\leq k\leq n$ one orbit of vertices $x\in X_{G}^{\leq n}$ with $h(x)=k$. Given a vertex $x$ with $h(x)=k$, there are only finitely many cubes in the sublevel set $X_{G}^{\leq n}$ that has $x$ as their bottom. It follows that there can only be finitely many orbits of cubes in $X_{G}^{\leq n}$. $\square$
Proposition 3.6.
For $x\leq y$ with $x\not\prec y$, $|(x,y)|$ is contractible.
Proof.
The proof uses Quillen’s conical contraction [41, Section 1.5] just as [14, Lemma 2.3] or [42, Lemma 4.4]. For that we need to define a function $c:(x,y)\to(x,y)$ such that there is $y_{0}$ in $(x,y)$ such that $z\geq c(z)\leq y_{0}$. In fact for any $z\in(x,y]$, let $c(z)$ be the largest element of $(x,z)$ satisfying $x\preceq c(z)$, and let $y_{0}=c(y)$. One checks that $c$ is well-defined and the condition $z\geq c(z)\leq y_{0}$ is satisfied for any $z\in(x,y]$. Therefore $|(x,y)|$ is contractible. $\square$
Proposition 3.7.
$X_{G}$ is homotopy equivalent to $|\mathcal{P}(G)|$, hence $X_{G}$ is also contractible.
Proof.
We again follow the standard proof here via building up $|\mathcal{P}(G)|$ from $X_{G}$ by attaching new subcomplexes in such a way that the homotopy type stays the same. Given a closed interval $[x,y]$, define $r([x,y])=h(y)-h(x)$. We attach the contractible complexes $|[x,y]|$ for $x\not\npreceq y$ to $X_{G}$ in increasing order of $r$ value, attaching $|[x,y]|$ along $|[x,y)\bigcup(x,y]|$ which is a suspension of $(x,y)$. But $(x,y)$ is contractible by Proposition 3.6. So the attaching process does not change the homotopy type of the complexes. And we have $X_{G}$ is homotopy equivalent to $|\mathcal{P}(G)|$. $\square$
Lemma 3.8.
Let $x\in\mathcal{P}(G)$ with $h(x)=m$, then the stabilizer $\mathrm{Stab}_{V(G)}(x)$ of $x$ in $V(G)$ is isomorphic to $G\wr\mathrm{S}_{n}$.
Proof.
Fix $\alpha$ with $[\alpha]=x$ and suppose $f\in V(G)$ fixes $x$. Then $[f\ast\alpha]=[\alpha]$ which means $\alpha^{-1}\ast f\ast\alpha=[1_{n},(\vec{g},\sigma),1_{n})]$ for some $(\vec{g},\sigma)\in G\wr S_{n}$. So we have a map $\mathrm{Stab}_{V(G)}(x)\mapsto G\wr S_{n}$. It has an inverse $(\vec{g},\sigma)\mapsto\alpha\ast[1_{n},(\vec{g},\sigma),1_{n})]\ast\alpha^{-1}$. So we have an isomorphism from $\mathrm{Stab}_{V(G)}(x)$ to $G\wr\mathrm{S}_{n}$. $\square$
To extend this to the cell stabilizers, let us make some notation first. Fix $n$ and let $J$ be a subset of $[n]=\{1,2,\cdots,n\}$, denote $F_{J}^{(n)}$ the elementary forest obtained from $1_{n}$ by attaching a caret at every $j\in J$. Furthermore, let $F_{j}^{(n)}:=F_{\{j\}}^{(n)}$ and $\lambda_{J}^{(n)}:=[F_{J}^{(n)},(\vec{1_{G}},\mathrm{id}),1_{n+|J|}]$.
Lemma 3.9.
Let $[x,y]$ be a cube in $X_{G}$ with $h(x)=n$. Then the stabilizer of $[x,y]$ is a finite index subgroup of $G\wr\mathrm{S}_{n}$. In particular, if $G$ is of type $F_{n}$, so does the stabilizer of $[x,y]$.
Proof.
First observe that an element of $V(G)$ fixes the cube $[x,y]$ if and only if it fixes both $x$ and $y$. Suppose $x=[\alpha]$, $h(x)=n$ and $f$ fixes $x$. By Lemma 3.8, we first have $\alpha^{-1}\ast f\ast\alpha\in G\wr\mathrm{S}_{n}$, where we identify $G\wr\mathrm{S}_{n}$ as a subgroup of $\mathscr{F}_{n,n}(G)$ as in the previous proof. Since $x\prec y$, there exist an elementary forest $F_{J}^{(n)}$ such that $y=[\alpha\ast\lambda_{J}^{(n)}]$, where $J$ is a subset of $[n]$. Now the element $f$ also fixes $y$, so $[f\ast\alpha\ast\lambda_{J}^{(n)}]=[\alpha\ast\lambda_{J}^{(n)}]$, which means $[\alpha^{-1}\ast f\ast\alpha\ast\lambda_{J}^{(n)}]=[\lambda_{J}^{(n)}]$. If we write $\alpha^{-1}\ast f\ast\alpha=(\vec{g},\sigma)$, to make $[\alpha^{-1}\ast f\ast\alpha\ast\lambda_{J}^{(n)}]=[\lambda_{J}^{(n)}]$, we just need $\sigma(J)=J$. So the stabilizer of $[x,y]$ can be identified with the subgroup $\{(\vec{g},\sigma)\in G\wr\mathrm{S}_{n}\mid\sigma(J)=J\}$ which is a finite index subgroup of $G\wr\mathrm{S}_{n}$. $\square$
3.4. Connectivity of the descending links
Let us give a description of the descending link first. Given a vertex $x$ in $\mathcal{P}(G)$. The descending star $\operatorname{{\operatorname{St}}{\downarrow}}$ of $x$ in $X_{G}$ consists of all the cubes of the form $[y,x]$ where $y\preceq x$. For such a cube $C=[y,x]$ let $\mathrm{bot}(C):=y$ be the map giving the bottom vertex. This is a bijection from the set of such cubes to the set $D(x):=\{y\in\mathcal{P}(G)\mid y\preceq x\}$. The cube $[y^{\prime},x]$ is a face of $[y,x]$ if and only if $y^{\prime}\in[y,x]$, if and only if $y^{\prime}\geq y$. Thus a cube $C^{\prime}$ is a face of $C$ if and only if $\mathrm{bot}(C^{\prime})\geq\mathrm{bot}(C)$ and $\mathrm{bot}$ is an order reversing map. Therefore, a simplex in $\operatorname{{\operatorname{Lk}}{\downarrow}}(x)$ can be identified with an element in $\mathcal{P}(G)$ such that $y\prec x$, where the dimension of the simplex plus one is just the number of elementary splits needed to get from $y$ to $x$. And face relation is reverse of the relation $<$ on $D(x)\setminus x$. Since $X_{G}$ is a cube complex, $\operatorname{{\operatorname{Lk}}{\downarrow}}(x)$ is a simplicial complex.
Now if $h(x)=n$, by left cancellation, $\operatorname{{\operatorname{Lk}}{\downarrow}}(x)$ is isomorphic to the following simplicial complex $\mathcal{E}_{n}(G)$: $[1_{n},(\vec{g},\sigma),F_{J}^{(n-|J|)}]$ for $J$ nonempty, and face relation given by reversing of the ordering $\leq$ in $\mathcal{P}(G)$.
Note that when $G$ is trivial, the simplices of $\mathcal{E}_{n}(1)$ are of the form $[1_{n},\sigma,F_{J}^{(n-|J|)}]$. But up to the equivalence relation on $\mathcal{F}_{1,n-|J|}(G)$ (i.e. right action by $S_{n-|J|}$), a $k$-simplex in $\mathcal{E}_{n}(1)$ is just a way to pick $(k+1)$ pair of roots in $1_{n}$. Hence it is isomorphic to the classical matching complex of $n$ points which is well-known to be highly connected. On the other hand, we have a forgetful map
$\pi:\mathcal{E}_{n}(G)\rightarrow\mathcal{E}_{n}(1)$
$[1_{n},(\vec{g},\sigma),F]\mapsto[1_{n},\sigma,F]$
We want to show that the map $\pi$ is a complete join so we can apply [26, Proposition 3.5] to show that $\mathcal{E}_{n}(G)$ is highly connected. Let us recall the definition here.
Definition 3.10.
A surjective simplicial map $\pi:Y\to X$ is called a complete join if it satisfies the following properties:
  1. (1)
    $\pi$ is injective on individual simplices.
  2. (2)
    For each $p$-simplex $\sigma=\langle v_{0},\cdots,v_{p}\rangle$ of $X$, $\pi^{-1}(\sigma)$ is the join $\pi^{-1}(v_{0})\ast\pi^{-1}(v_{1})\ast\cdots\ast\pi^{-1}(v_{p})$.
Lemma 3.11.
Let $[1_{n},(\vec{g}_{1},\sigma_{1}),F_{i}^{(n-1)}]$ and $[1_{n},(\vec{g}_{2},\sigma_{2}),F_{J}^{(n-k-1)}]$ be two simplicies of dimension 00 and $k$ in $\mathcal{E}_{n}(G)$. Suppose that their images under the map $\pi$ are contained in a simplex of $\mathcal{E}_{n}(1)$. Then there exists a simplex in $\mathcal{E}_{n}(G)$ containing both of them.
Proof.
We may assume that $[1_{n},(\vec{g}_{1},\sigma_{1}),F_{i}^{(n-1)}]$ is not a face of $[1_{n},(\vec{g}_{2},\sigma_{2}),F_{J}^{(n-k-1)}]$. There is an action of $G\wr\mathrm{S}_{n}$ on $\mathcal{E}_{n}(G)$ given by
$(\vec{g}^{\prime},\sigma^{\prime})[1_{n},(\vec{g},\sigma),E]=[1_{n},(\vec{g}^{\prime},\sigma^{\prime})(\vec{g},\sigma),E]$
One checks easily that for each $k\geq 0$, the action is transitive on the $k$-simplices of $\mathcal{E}(G)$. We can therefore assume without loss of generality that $[1_{n},(\vec{g}_{2},\sigma_{2}),F_{J}^{(n-k-1)}]=[1_{n},(\vec{1}_{G},\mathrm{id}),F_{[k]}^{(n-k-1)}]$ where $F_{[k]}^{(n-k-1)}$ is the forest with the first $k+1$ components being carets and the other $n-2k-2$ components all single roots.
Since the image of $[1_{n},(\vec{g}_{1},\sigma_{1}),F_{i}^{(n-1)}]$ and $[1_{n},(\vec{1}_{G},\mathrm{id}),F_{[k]}^{(n-k-1)}]$ under $\pi$ are different, $\sigma^{-1}_{1}(i),\sigma^{-1}_{1}(i+1)$ must be disjoint from $\{1,2,\cdots,2k+2\}$. Up to equivalence (i.e. right action of $G\wr S_{n-1}$) we can also assume that $i,i+1$ are disjoint from $\{1,2,\cdots,2k+2\}$ and the first $22k+$ 2 coordinates of $\vec{g}_{1}$ are $1_{G}$, in particular we can assume that $\sigma_{1}$ leaves $\{1,2,\cdots,2k+2\}$ pointwisely fixed. Now it is easy to see that $[1_{n},(\vec{g}_{1},\sigma_{1}),F_{[k]}^{(n-k-1)}]=[1_{n},(\vec{1}_{G},\mathrm{id}),F_{[k]}^{(n-k-1)}]$, so $[1_{n},(\vec{g}_{1},\sigma_{1}),F_{[k]\cup\{i\}}^{(n-k-2)}]$ is a simplex containing $[1_{n},(\vec{1}_{G},\mathrm{id}),F_{[k]}^{(n-k-1)}]$ and $[1_{n},(\vec{g}_{1},\sigma_{1}),F_{i}^{(n-1)}]$ as faces. $\square$
Lemma 3.12.
$\pi$ is a complete join.
Proof.
By definition it is easy to see that the map $\pi$ is surjective and injective on individual simplices. Let $e$ be a $k$-simplex in $\mathcal{E}(1)$ with vertices $v_{0},\ldots,v_{k}$. To prove $\pi$ is a complete join, it just remains to show
$\pi^{-1}(e)=\underset{0\leq j\leq k}{*}\pi^{-1}(v_{j})$
$\subseteq$ ”: This inclusion just says that vertices in $\pi^{-1}(\sigma)$ that are connected by an edge map to distinct vertices under $\pi$ which is clear.
$\supseteq$ ”: We prove this by induction on $k$. Suppose $\pi^{-1}(e)\supseteq\underset{0\leq j\leq k}{*}\pi^{-1}(v_{j})$ for $k=r$. Now given an $(r+1)$-simplex $e^{\prime}=\langle v_{0},\cdots,v_{r+1}\rangle$ which is a join of $e=\langle v_{0},\cdots,v_{r}\rangle$ and $v_{r+1}$. We just need to show for any simplex $\bar{e},\bar{v}_{r+1}\in\mathcal{E}(G)$ such that $\pi(\bar{e})=e,\pi(\bar{v}_{r+1})=v_{r+1}$, we have a $(r+1)$-simplex contains both $\bar{e}$ and $v_{r+1}$. But follows immediately from Lemma 3.11 so we complete the proof. $\square$
Corollary 3.13.
$\mathcal{E}(G)$ is $(\lfloor\frac{n+1}{3}\rfloor-2)$connected.
Proof.
The matching complex $\mathcal{M}(K_{n})$ is $(\lfloor\frac{n+1}{3}\rfloor-2)$-connected, see for example [14, Proposition 3.16]. And observe that the link of a $k$-simplex is isomorphic to $\mathcal{M}(K_{n-2k})$ , which is $(\lfloor\frac{n-2k+1}{3}\rfloor-2)$connected. Hence $\mathcal{M}(K_{n})$ is wCM of dimension $(\lfloor\frac{n+1}{3}\rfloor-1)$. Recall that a simplicial complex wCM of dimension $n$ if it is $(n-1)$-connected and the link of each $k$-simplex is $(n-k-2)$-connected. Now by [26, Proposition 3.5], $\mathcal{E}(G)$ is also wCM of dimension $(\lfloor\frac{n+1}{3}\rfloor-1)$, in particular, it’s $(\lfloor\frac{n+1}{3}\rfloor-2)$connected. $\square$
Theorem 3.14.
$V(G)$ is of type $F_{m}$ if and only if $G$ is of type $F_{m}$.
Proof.
By Corollary 1.14, it suffices to prove the “if” direction. We apply 3.2. Assume $G$ is of type $F_{m}$, consider then action $V(G)$ on the cubicial complex $|S_{G}|$ with the height function $h$. By Proposition 3.5, the action of $V(G)$ on $X_{G}$ is cocompact; by Lemma 3.9, the stabilizer of a cell $[z,y]$ is of type $F_{m}$ ; and by Corollary 3.13 the high connectivity of the descending link of a vertex is guaranteed when its height is big enough. So by 3.2, $V(G)$ is of type $F_{m}$. $\square$
Remark 3.15.
One can also define the labeled Thompson group $T(G)$ corresponding to the Thompson group $T$. As a subgroup of $V(G)$, it is the preimage of $T$ under the canonical projection $V(G)\to V$. Our proof for the finiteness property of $V(G)$ works almost without change for $T(G)$. The only place one needs a little extra care is in the proof of Corollary 3.13, we use the fact the matching complex over the complete graph is highly connected. For $T(G)$, one has to use the matching complex over the cyclic graph which is also known to be highly connected, see for example [48, Lemma 3.19]. Therefore, we have $T(G)$ is of type $F_{n}$ if and only if $G$ is.
4. Tools to compute the bounded cohomology of groups
We collect some tools that can be used to show a group is boundedly acyclic in this section. The readers should consult [22] for the basics of bounded cohomology. The following cohomological techniques are developed in [35] to compute the bounded cohomology of groups.
Definition 4.1.
Let $X$ be a set. We call a binary relation $\perp$ generic if for every given finite set $Y\subseteq X$, there is an element $x\in X$ such that $y\perp x$ for all $y\in Y$.
Recall that any generic relation on $X$ gives rise to a semisimplicial set $X_{\bullet}^{\perp}$ in the following way: we define $X_{n}^{\perp}$ to be the set of all $(n+1)$-tuples $\left(x_{0},\ldots,x_{n}\right)\in X^{n+1}$ for which $x_{i}\perp x_{j}$ holds for all $0\leq j<i\leq n$. The face maps $\partial_{n}:X_{n}^{\perp}\rightarrow X_{n-1}^{\perp}$ are the usual simplex face maps, i.e.,
$\partial_{n}\left(x_{0},\ldots,x_{n}\right)=\sum_{i=0}^{n}(-1)^{i}\left(x_{0},\ldots,\widehat{x}_{i},\ldots,x_{n}\right),$
where $\widehat{x}_{i}$ means that $x_{i}$ is omitted.
Lemma 4.2.
[35, Proposition 3.2] If the binary relation $\perp$ is generic, then the complex $X_{\bullet}^{\perp}$ is boundedly acyclic (and connected).
Proposition 4.3.
[35, Theorem 3.3 + Remark 3.4] Let $\Gamma$ be a group acting on a boundedly acyclic connected semisimplicial set $X_{\bullet}$. Suppose further that:
  1. (i)
    The stabilizers of any point in $X_{p}$ are boundedly acyclic.
  2. (ii)
    The $\Gamma$-action on $X_{p}$ has finitely many orbits.
Then there is an isomorphism $H_{b}^{p}(\Gamma,\mathbb{R})\cong H_{b}^{p}\left(X_{\bullet}/\Gamma,\mathbb{R}\right)$ for every $p\geq 0$.
Combining Lemma 4.2 and Proposition 4.3, we have the following corollary.
Corollary 4.4.
Let $\Gamma$ be a group acting transitively on a set $X$ on which there is a generic relation $\perp$ preserved by $\Gamma$. If for every $n\geq 1$, the action of $\Gamma$ on $X^{\perp}_{n}$ is transitive and the stabilizers are boundedly acyclic, then $\Gamma$ is boundedly acyclic.
We also need the following criterion to prove the boundedly acyclicity for stabilizers.
Theorem 4.5.
[34, Corollary 5] Let $\Gamma$ be a group and $\Gamma_{0}<\Gamma$ a subgroup with the following two properties:
(i) every finite subset of $\Gamma$ is contained in some $\Gamma$-conjugate of $\Gamma_{0}$,
(ii) $\Gamma$ contains an element $g$ such that the conjugates of $\Gamma_{0}$ by $g^{p}$ and by $g^{q}$ commute for all $p\neq q$ in $\mathbb{Z}$.
Then $\mathrm{H}_{\mathrm{b}}^{n}(\Gamma,\mathbb{R})=0$ holds for all $n>0$.
We also use the following powerful generalization of 4.5 for proving bounded acyclicity.
Definition 4.6.
[16, Definition 1.1] A group $\Gamma$ has commuting cyclic conjugates if for every finitely generated subgroup $H\leq\Gamma$ there exist $t\in\Gamma$ and $k\in\mathbb{N}_{\geq 2}\cup\{\infty\}$ such that:
(1) $\left[H,t^{p}Ht^{-p}\right]=1$ for all $1\leq p<k$;
(2) $\left[H,t^{k}\right]=1$.
Here we read that $t^{\infty}=1$.
Theorem 4.7.
[16, Theorem 1.3] Let $\Gamma$ be a group with commuting cyclic conjugates. Then $\Gamma$ is boundedly acyclic.
Recall that a subgroup $H\leq\Gamma$ is co-amenable in $\Gamma$ if there is a $\Gamma$-invariant mean on $\Gamma/H$. Since Γ/[$\Gamma/[\Gamma,\Gamma]$ is abelian hence amenable, $[\Gamma,\Gamma]$ is co-amenable in $\Gamma$.
Lemma 4.8.
[33, Proposition 8.6.6] Let $\Gamma$ be a group, $\Gamma_{0}<\Gamma$ a co-amenable subgroup. Then the restriction map
$\mathrm{H}_{\mathrm{b}}^{\bullet}(\Gamma,\mathbb{R})\longrightarrow\mathrm{H}_{\mathrm{b}}^{\bullet}\left(\Gamma_{0},\mathbb{R}\right)$
is injective.
5. Bounded acyclicity of $\phi$-labeled Thompson groups
In this section we prove that the labeled Thompson groups are all boundedly acyclic based on ideas in [2, 34, 35].
Given a group $G$ with injective wreath recursion $\phi$, denote the corresponding labeled Thompson group by $V_{\phi}(G)$. A dyadic interval is a subset of =$\mathfrak{C}=\{0,1\}^{\omega}$ of the form $u\{0,1\}^{\omega}$ for some $u\in\{0,1\}^{<\omega}$. Recall that dyadic intervals are clopen and the set of dyadic intervals form a topological basis for $\mathfrak{C}$. A dyadic neighborhood of $x\in\mathfrak{C}$ is a dyadic interval containing $x$.
Definition 5.1.
Given an element $\alpha\in V_{\phi}(G)$, with $\alpha=\scriptsize\begin{bmatrix}u&...\\ g&...\\ v&...\end{bmatrix}$, then we say $\scriptsize\begin{bmatrix}u\\ g\\ v\end{bmatrix}$ is a labeled arrow of $\alpha$ at $U:=u\{0,1\}^{\omega}$. Also we define its label at $U$ to be $g$, denoted by $l(\alpha|_{U})=g$, where $g\in G$.
Note that the label depends on the choice of dyadic intervals and even the label at a dyadic interval is not $1_{G}$, when we restrict to a subinterval it might be $1_{G}$. Therefore the label at a point is not well-defined unless the label is $1_{G}$ at some neighborhood of this point.
Example 5.2.
Consider the right wreath recursion: $\phi_{r}:g\rightarrow{((1,g),id_{S_{2}})}$. Given $g_{0}\in G$, let $\Lambda_{u}(g_{0})=\scriptsize\begin{bmatrix}u&u_{2}&\cdots&u_{n}\\ g_{0}&1&\cdots&1\\ u&u_{2}&\cdots&u_{n}\end{bmatrix}$ where $\{u,u_{2},\cdots,u_{n}\}$ is a partition set. Then restricted to any dyadic interval not containing the point ${u}{11}\cdots$, its label is $1_{G}$.
Recall from Remark 1.7 that $\alpha=\scriptsize\begin{bmatrix}u&...\\ g&...\\ v&...\end{bmatrix}$ acts on $\{0,1\}^{\omega}$ by mapping $uw$ to $v(w)\phi^{*}(g)$ for all $w\in\{0,1\}^{\omega}$, and we use $(w)\alpha$ to denote the image of $w$ under the action of $\alpha$. In particular the image $U\alpha$ is again a dyadic interval. Given two elements $\alpha,\beta$, if $l(\alpha|_{U})=g$ and $l(\beta|_{U\alpha})=h$, we have $l(\alpha\beta|_{U})=gh$.
Definition 5.3.
The labeled support of $\alpha\in V_{\phi}(G)$ denoted by $\mathrm{lsupp}(\alpha)$ is defined as the complement set of the union of all dyadic neighborhoods $U$ satisfying $l(\alpha|_{U})=1_{G}$ and the action of $\alpha$ restricting to $U$ is the identity map.
Note that the labeled support of an element depends strongly on the wreath recursion.
Example 5.4.
Again consider $\Lambda_{u}(g_{0})=\scriptsize\begin{bmatrix}u&u_{2}&\cdots&u_{n}\\ g_{0}&1&\cdots&1\\ u&u_{2}&\cdots&u_{n}\end{bmatrix}$. Then for the diagonal wreath recursion, $\Delta:g\rightarrow{((g,g),id_{S_{2}})}$, $\mathrm{lsupp}(\Lambda_{u}(g_{0}))=u\{0,1\}^{\omega}$. But for the right diagonal wreath recursion, $\phi_{r}:g\rightarrow{((1,g),id_{S_{2}})}$, $\mathrm{lsupp}(\Lambda_{u}(g_{0}))=\{{u}{11}\cdots\}$ which is closed but not open.
The labeled support of a subset of $V_{\phi}(G)$ is defined as the union of the labeled support of all elements in this set.
Lemma 5.5.
Two elements with disjoint labeled supports commute.
Proof.
Given two elements $\alpha,\beta$ with disjoint labeled supports, then after some expansions they can be written as
$\alpha=\scriptsize\begin{bmatrix}u_{1}&\cdots&u_{k}&v_{1}&\cdots&v_{k^{\prime}}\\ 1_{G}&\cdots&1_{G}&g_{1}&\cdots&g_{k^{\prime}}\\ u_{1}&\cdots&u_{k}&w_{1}&\cdots&w_{k^{\prime}}\end{bmatrix},\beta=\scriptsize\begin{bmatrix}w_{1}&\cdots&w_{k^{\prime}}&u_{1}&\cdots&u_{k}\\ 1_{G}&\cdots&1_{G}&h_{1}&\cdots&h_{k}\\ w_{1}&\cdots&w_{k^{\prime}}&t_{1}&\cdots&t_{k}\end{bmatrix},$
where these $u_{i}\{0,1\}^{\omega}$ together with $v_{j}\{0,1\}^{\omega}$ form a dyadic partition of $\mathfrak{C}$.
So
$\alpha\beta=\scriptsize\begin{bmatrix}u_{1}&\cdots&u_{k}&v_{1}&\cdots&v_{k^{\prime}}\\ h_{1}&\cdots&h_{k}&g_{1}&\cdots&g_{k^{\prime}}\\ t_{1}&\cdots&t_{k}&w_{1}&\cdots&w_{k^{\prime}}\par\end{bmatrix}.$
Note we can also write
$\alpha=\scriptsize\begin{bmatrix}t_{1}&\cdots&t_{k}&v_{1}&\cdots&v_{k^{\prime}}\\ 1_{G}&\cdots&1_{G}&g_{1}&\cdots&g_{k^{\prime}}\\ t_{1}&\cdots&t_{k}&w_{1}&\cdots&w_{k^{\prime}}\end{bmatrix},\beta=\scriptsize\begin{bmatrix}v_{1}&\cdots&v_{k^{\prime}}&u_{1}&\cdots&u_{k}\\ 1_{G}&\cdots&1_{G}&h_{1}&\cdots&h_{k}\\ v_{1}&\cdots&v_{k^{\prime}}&t_{1}&\cdots&t_{k}\end{bmatrix}.$
One checks easily now that $\alpha\beta=\beta\alpha$. $\square$
Definition 5.6.
Let $\omega_{0}:=000\cdots$ be the constant 00 word in $\mathfrak{C}$. Two elements $\alpha_{1},\alpha_{2}\in V_{\phi}(G)$ are called equivalent if there is a dyadic neighborhood $U$ of $\omega_{0}$ such that if $(w)\alpha_{1}=(w)\alpha_{2}$ for all $w\in U$ and $l(\alpha_{1}|_{U})=l(\alpha_{2}|_{U})$. A labeled germ is an equivalence class under this relation. Given $\alpha\in V_{\phi}(G)$, we use $[\alpha]$ to denote its equivalence class. Denote the set of labeled germs by $\mathcal{G}$.
Define a $V_{\phi}(G)$ right action on $\mathcal{G}$ by $[\alpha]\beta:=[\alpha\beta]$. This is a well-defined action: with the same notation, given $\alpha_{1}$ equivalent to $\alpha_{2}$, the corresponding dyadic neighborhood $U$ small enough so that the label of $\beta$ on $(U)\alpha$ is also well-defined, then $(w)\alpha_{1}\beta=(w)\alpha_{2}\beta$ for all $w\in U$ and $l(\alpha_{1}\beta|_{U})=l(\alpha_{1}|_{U})l(\beta|_{U\alpha_{1}})=l(\alpha_{2}|_{U})l(\beta|_{U\alpha_{2}})=l(\alpha_{2}\beta|_{U})$.
We proceed to define a binary relation $\perp$ on $\mathcal{G}$.
Definition 5.7.
Given two labeled germs $[\alpha_{1}],[\alpha_{2}]$, we say $[\alpha_{1}]\perp[\alpha_{2}]$ if $(\omega_{0})\alpha_{1}\neq(\omega_{0})\alpha_{2}$.
Remark 5.8.
If $(\omega_{0})\alpha_{1}\neq(\omega_{0})\alpha_{2}$, we obviously have $[\alpha_{1}]\neq[\alpha_{2}]$. But the converse is not necessarily true.
Lemma 5.9.
$\perp$” defines a generic relation on $\mathcal{G}$ and $V_{\phi}(G)$ preserves the relation.
Proof.
The action of $V_{\phi}(G)$ obviously preserves the relation. To show that $\perp$” defines a generic relation on $\mathcal{G}$ it suffices to observe that the orbit of the element $\omega_{0}$ under the action of $V_{\phi}(G)$ on $\{0,1\}^{\omega}$ has infinitely many elements. $\square$
Let $\mathcal{G}^{\perp}_{n}$ denote the set of generic $(n+1)$-tuples of $\mathcal{G}$. We first note the following lemma, where $V$ is the Thompson group $V$. Not that when $\gamma\in V$ is regarded as an element of $V_{\phi}(G)$ with trivial labels, then its labeled support equals its support as a homeomorphism on the Cantor set.
Lemma 5.10.
[8, Section 1.1] For any clopen set $A,B,C$ in the Cantor set $\mathfrak{C}$ with $B,C$ properly contained in $A$, there exists a $\gamma\in V$ such that $\mathrm{supp}(\gamma)\subset A$ and $B\gamma\subset C$.
Lemma 5.11.
Given clopen sets $A,B$ in $\mathfrak{C}$ such that $A$ is properly contained in $B$, then there exist a $\gamma\in V$ with support in $B$ such that $A\gamma^{k}\cap A=\varnothing$ for every $k>0$
Proof.
Viewing $\mathfrak{C}$ as the subset of the real interval $[0,1]$ under the standrad definition, if $[0,\frac{1}{2}]\cap\mathfrak{C}\subset B$ and $A\subset[\frac{3}{16},\frac{1}{4}]$, then we pick an element $\lambda$ in the Thompson group $F$ such that $\lambda(x)=\frac{x}{2}$ when $x\in[0,\frac{1}{4}]\cap\mathfrak{C}$ and $\lambda$ is identity outside $[0,\frac{1}{2}]$. Hence we are done. If not, then by Lemma 5.10 we can find a $\tau$ such that $[0,\frac{1}{2}]\cap\mathfrak{C}\subset\tau(B)$ and $\tau(A)\subset[\frac{3}{16},\frac{1}{4}]$, now $\gamma^{\tau}$ is an element in $V$ fulfills the conditions in the lemma. $\square$
Lemma 5.12.
$V_{\phi}(G)$ acts transitively on $\mathcal{G}^{\perp}_{n}$ for all $n\geq 1$.
Proof.
Given two generic tuples $([\alpha_{0}],\cdots,[\alpha_{n}])$, $([\beta_{0}],\cdots,[\beta_{n}])\in\mathcal{G}^{\perp}_{n}$, we want to find an element in $V_{\phi}(G)$ that maps $[\alpha_{i}]$ to $[\beta_{i}]$. Note first that $(\omega_{0})\alpha_{0},\cdots,(\omega_{0})\alpha_{n}$ are pairwise distinct by definition, and the same holds for $(\omega_{0})\beta_{0},\cdots,(\omega_{0})\beta_{n}$. Choose a small dyadic neighborhood $U$ of $\omega_{0}$ such that both finite sets $\{U\alpha_{0},\cdots,U\alpha_{n}\}$ and $\{U\beta_{0},\cdots,U\beta_{n}\}$ are pairwise disjoint dyadic intervals. Then there is an element $\gamma\in V_{\phi}(G)$ such that $l(\gamma|_{U\beta_{i}})=l(\beta_{i}|_{U})^{-1}l(\alpha_{i}|_{U})$ and $(w)\beta_{i}\gamma=(w)\alpha_{i}$ for all $w\in U$. In more detail, we can find two dyadic partitions $\mathcal{U}=\{u_{1},\cdots,u_{k}\},\mathcal{W}=\{v_{1},\cdots,v_{k}\}$ for some $k\geq n$ such that $u_{i}\{0,1\}^{\omega}=U\alpha_{i}$ and $v_{i}\{0,1\}^{\omega}=U\beta_{i}$ for $1\leq i\leq n$. Define now $\gamma\in V$ by the $G$-matrix $\scriptsize\begin{bmatrix}v_{1}&\cdots&v_{n}&v_{n+1}&\cdots&v_{k}\\ h_{1}&\cdots&h_{n}&1_{G}&\cdots&1_{G}\\ u_{1}&\cdots&u_{n}&u_{n+1}&\cdots&u_{k}\end{bmatrix}$. where $h_{i}:=l(\beta_{i}|_{U})^{-1}l(\alpha_{i}|_{U})\in G$ for $1\leq i\leq n$. Now it’s easy to check that $\gamma$ maps each $[\alpha_{i}]$ to $[\beta_{i}]$. $\square$
Lemma 5.13.
The stabilizers for the $V_{\phi}(G)$-action on $\mathcal{G}^{\perp}_{n}$ are boundedly acyclic for all $n\geq 0$.
Proof.
Let $([x_{0}],\cdots[x_{n}])\in\mathcal{G}^{\perp}_{n}$ where each $(\omega_{0})x_{i}$ belongs to $00\{0,1\}^{\omega}$ for all $0\leq i\leq n$. Since the action of $V_{\phi}(G)$ on $\mathcal{G}^{\perp}_{n}$ is transitive, it suffices to prove that the stabilizer $G_{n}$ of $([x_{0}],\cdots[x_{n}])$ is boundedly acyclic. By definition, it’s easy to see that $G_{e}=\{\gamma\in V_{\phi}(G):\omega_{0}\notin\mathrm{lsupp}(\gamma)\}$ where $G_{e}=\mathrm{Stab}_{V_{\phi}(G)}([e])$ and $e$ is the identity element in $V(G)$. Hence we have $G_{n}=\bigcap_{i=0}^{n}G_{e}^{x_{i}}=\{\gamma\in V_{\phi}(G):(\omega_{0})x_{i}\notin\mathrm{lsupp}(\gamma)\text{ for each }0\leq i\leq n\}$. Now let $\Omega=10\{0,1\}^{\omega}$, denote by $G_{\Omega}$ the subgroup consisting of elements with labeled support in $\Omega$, so $G_{\Omega}\leq G_{n}$. We claim the pair $(G_{n},G_{\Omega})$ satisfies the conditions in 4.5, hence $G_{n}$ is boundedly acyclic.
First we claim every finite subset F of $G_{n}$ can be conjugated by an element of $V\cap G_{n}$ into $G_{\Omega}$. In fact, $\mathrm{lsupp}(F)$ is a closed hence proper subset of $\mathfrak{C}-\{x_{0}(\omega_{0}),\cdots,x_{n}(\omega_{0})\}$, so we could choose clopen subsets $A,B$ of $\mathfrak{C}-\{x_{0}(\omega_{0}),\cdots,x_{n}(\omega_{0})$ such that $\mathrm{lsupp}(F)\subsetneq B\subsetneq A$ and $\Omega\subsetneq A$. Now by Lemma 5.10 we can pick a $\gamma_{1}\in V\cap G_{n}$ such that $B\gamma_{1}\subset\Omega$, hence $\mathrm{lsupp}(F)\gamma_{1}\subset\Omega$ and $F^{\gamma_{1}}\subset G_{\Omega}$.
Next, we define an element $\gamma_{2}\in V$ as follows:
$\gamma_{2}(uw)=\begin{cases}00w&\text{ if }u=00\\ 010w&\text{ if }u=01\\ 011w&\text{ if }u=10\\ 1w&\text{ if }u=11\\ \end{cases}$
Note that $\gamma_{2}\in V\cap G_{n}$ and $\Omega\gamma_{2}^{k}\cap\Omega=\varnothing$ for all $k\in\mathbb{Z}-\{0\}$. Therefore by Lemma 5.5 the conjugation subgroup of $G_{\Omega}$ by $\gamma_{2}^{p}$ and $\gamma_{2}^{q}$ always commute if $p\neq q$. $\square$
Remark 5.14.
One could also use 4.7 to prove the Lemma.
Now combining Lemma 5.12 and 5.13 with Corollary 4.4 we have:
Theorem 5.15.
$V_{\phi}(G)$ is boundedly acyclic for any injective wreath recursion $\phi$.
6. Bounded cohomology of some topological full groups
In this section, we prove topological full groups with extremely proximal actions are boundedly acyclic. Recall given a subgroup $\Gamma\subseteq\operatorname{Homeo}(\mathfrak{C})$, its topological full group is defined as:
$\llbracket\Gamma\rrbracket=\left\{g\in\operatorname{Homeo}(\mathfrak{C}):\begin{array}[]{l}\text{ for each }x\in\mathfrak{C}\text{ there exists a neighbour- }\\ \text{ hood }U\text{ of }x\text{ and }\gamma\in\Gamma\text{ such that }\left.g\right|_{U}=\left.\gamma\right|_{U}\end{array}\right\}$
Note by definition, we have $\llbracket\llbracket\Gamma\rrbracket\rrbracket=\llbracket\Gamma\rrbracket$ and a subgroup $\Gamma\subset\operatorname{Homeo}(\mathfrak{C})$ is a topological full group if and only if $\llbracket\Gamma\rrbracket=\Gamma$.
Definition 6.1.
Let $\Gamma$ be a group acts on the Cantor set, the action is extremely proximal if for any nonempty and proper clopen sets $V_{1},V_{2}\subsetneq\mathfrak{C}$ there exists $\gamma\in\Gamma$ such that $\gamma(V_{2})\subsetneq V_{1}$.
Inspired by the uniform simplicity results of topological full groups in [23, Section 5], in this section we will prove the following:
Theorem 6.2.
Let $\Gamma$ be a topological full group such that the action is extremely proximal. Then $\Gamma$ is boundedly acyclic.
Proof.
First let $p\in\mathfrak{C}$, we define $X_{\Gamma}:=\Gamma/\sim$ where $\gamma_{1}\sim\gamma_{2}$ when they agree on a neighborhood of $p$, let $[\gamma]\in X_{\Gamma}$ denote the equivalence class of $\gamma$. View $X_{\Gamma}$ as a $\Gamma$-set where $\Gamma$ acts by $[\gamma_{1}]\gamma_{2}:=[\gamma_{1}\gamma_{2}]$. Then we define a relation in $X_{\Gamma}$ by setting $[\gamma_{1}]\perp[\gamma_{2}]$ if $(p)\gamma_{1}\neq(p)\gamma_{2}$. Now it’s easy to see that the right $\Gamma$ action preserves the relation $\perp$ and $\perp$ is actually generic in the sense of Definition 4.1. So we get a boundedly acyclic $\Gamma$-semi-simplicial set $(X_{\Gamma})_{\bullet}^{\perp}$ by Lemma 4.2. For disjoint clopen sets $A,B$ and $\tau\in\Gamma$ maping $A$ to $B$ homeomorphically, define a homeomorphism
$\gamma_{\tau,A,B}(x):=\begin{cases}\tau(x)&x\in A\\ \tau^{-1}(x)&x\in B\\ x&otherwise\end{cases}$
By definition $\gamma_{\tau,A,B}\in\Gamma$ and is of order $2$.
We want to apply Corollary 4.4 to show $\Gamma$ is boundedly acyclic. For that we first prove that for each $n\geq 0$ the $\Gamma$ action on the set $(X_{\Gamma})_{n}^{\perp}$ is transitive. Let us prove it by induction on $n$. For $n=0$ it is trivial, suppose that the actions are transitive in $<n$ cases. Given $([\gamma_{0}],\cdots,[\gamma_{n}]),([\gamma^{\prime}_{0}],\cdots,[\gamma^{\prime}_{n}])\in(X_{\Gamma})_{n}^{\perp}$, by the induction hypothesis there is $\gamma\in\Gamma$ with $[\gamma_{i}]\gamma=[\gamma^{\prime}_{i}]$ for $i<n$. We first claim that $\gamma$ can be chosen such that $(p)\gamma_{n}\gamma\neq(p)\gamma_{n}^{\prime}$. In fact, if $(p)\gamma_{n}\gamma=(p)\gamma_{n}^{\prime}$, then we can modify $\gamma$ in the following way. Choose $A_{1}$ to be a small clopen subset disjoint from $(p)\gamma_{0}\gamma,\cdots,(p)\gamma_{n-1}\gamma$ but containing $(p)\gamma_{n}\gamma$, and $B_{1}$ to be another clopen subset disjoint from $(p)\gamma_{0}\gamma,\cdots,(p)\gamma_{n-1}\gamma$ and $A_{1}$. If we replace $\gamma$ with $\gamma\gamma_{\tau_{1},A_{1},B_{1}}$, then $(p)\gamma_{n}\gamma\neq(p)\gamma_{n}^{\prime}$ and $[\gamma_{i}]\gamma=[\gamma^{\prime}_{i}]$ for $i<n$. Now let $\alpha:=\gamma_{n}\gamma$, $\beta:=\alpha^{-1}\gamma^{\prime}_{n}$ , they are elements of $\Gamma$. Note that $(p)\alpha\neq(p)\alpha\beta$. Pick a small clopen set $A_{2}$ containing $(p)\alpha$ but disjoint from $(p)\gamma^{\prime}_{0},\cdots,(p)\gamma^{\prime}_{n-1}$, such that its image $B_{2}=A_{2}\beta$ is also disjoint from $(p)\gamma^{\prime}_{0},\cdots,(p)\gamma^{\prime}_{n-1}$. Note now that the element $\gamma_{\beta,A_{2},B_{2}}\in\Gamma$, fixes $\{(p)\gamma^{\prime}_{0},\cdots,(p)\gamma^{\prime}_{n-1}\}$ pointwise and maps $(p)\alpha$ to $(p)\alpha\beta$. One checks also that $([\gamma_{0}],\cdots,[\gamma_{n}])\gamma\gamma_{\beta,A_{2},B_{2}}=([\gamma^{\prime}_{0}],\cdots,[\gamma^{\prime}_{n}])$.
Next we prove that for each $n\geq 0$ the stabilizers of $\Gamma$-action on $(X_{\Gamma})_{n}^{\perp}$ is boundedly acyclic. Let $([\gamma_{0}],\cdots,[\gamma_{n}])\in(X_{\Gamma})_{n}^{\perp}$ and $\Gamma_{n}$ be its stabilizer, let $F:=\{(p)\gamma_{0},\cdots,(p)\gamma_{n}\}$. Then $\Gamma_{n}$ consists of homeomorphisms that fix a neighborhood of $F$ pointwisely. We will prove that $\Gamma_{n}$ is boundedly acyclic by showing that it has commuting cyclic conjugates (see Definition 4.6). For a given finitely generated group $H\leq\Gamma_{n}$, there is a neighborhood $U$ of $F$ that is fixed by every element of $H$ pointwisely. Pick a clopen set $U^{\prime}$ with $F\subsetneq U^{\prime}\subsetneq U$. Since the $\Gamma$-action on $\mathfrak{C}$ is extremely proximal, there is $\omega\in\Gamma$ such that $(\mathfrak{C}-U)\omega\subset U-U^{\prime}$. Note that $\gamma_{\omega,\mathfrak{C}-U,(\mathfrak{C}-U)\omega}\in\Gamma_{n}$ and is of order 2, and $H$ commutes with its conjugate $H^{\gamma_{\omega,\mathfrak{C}-U,(\mathfrak{C}-U)\omega}}$ since they have disjoint support. Therefore $\Gamma_{n}$ has the commuting cyclic conjugates for $k=2$ in Definition 4.6. Hence by 4.7, $\Gamma_{n}$ is boundedly acyclic. This finishes the proof. $\square$
In [23, Section 5], Gal and Gismatullin proved that topological full groups acting on $\mathfrak{C}$ extremely proximally are uniformly simply. Recall a group is uniformly simple if for any given conjugacy class $C$, every element of the group can be written as the product of bounded number of elements from $C^{\pm}$. They checked the commutator subgroup of the Neretin groups and the commutator subgroups of the Higman–Thompson groups satisfy these assumptions, hence they are also boundedly acyclic by our theorem. But a group is always coamenable with its commutator subgroup, hence we have the following corollary by Lemma 4.8.
Corollary 6.3.
Neretin groups (viewed as discrete groups) and Higman–Thompson groups are boundedly acyclic.
It is also the case that Röver–-Nekrashevych groups are topological full groups (see for example [3, Example 5.18]) and act on the Cantor set extremely proximally since the Higman–Thompson groups already do.
Corollary 6.4.
The Röver–Nekrashevych groups are all boundedly acyclic.
Another class of groups satisfies the assumptions of our theorem is the twisted Brin–Thompson groups of Belk and Zaremsky [6]. In fact, it is known that they are topological full group, and act extremely proximal on the Cantor space, see the proof of [3, Corollary 5.14, Proposition 5.23].
Corollary 6.5.
The twisted Brin Thompson groups are all boundedly acyclic.
In general, topological full groups could have non-trivial bounded cohomology.
Example 6.6.
Let $G$ be a countable group which acts on the Cantor set $\{0,1\}^{G}$ by permuting the coordinates using right multiplication. Then the corresponding topological full group $\llbracket G\rrbracket$ retracts to $G$. In fact let $w_{0}=\cdots 000\cdots$ be the constant 00 sequence in $\{0,1\}^{G}$, then it is a fixed point of $G$ hence also $\llbracket G\rrbracket$. Note that the clopen subsets of the form $\{w\in\{0,1\}^{G}\mid w_{g}=a_{g},g\in F\}$ where $F$ is a finite subset of $G$ and $a_{g}\in\{0,1\}$ form a topological basis of $\{0,1\}^{G}$. In particular, given an element of $\llbracket G\rrbracket\to G$, one can find such a clopen neighborhood $U_{0}$ of $w_{0}$, such that the action restricted to $U_{0}$ determines an element of $G$. One further checks that this element does not depend on the choice of $U_{0}$. Thus we have a well-defined retraction map $r:\llbracket G\rrbracket\to G$. In fact, let $\iota$ be the natural inclusion map from $G$ to $\llbracket G\rrbracket$, we have $r\circ\iota=\mathrm{id}_{G}$. Hence $r$ induces a splitting injective map on bounded cohomology. In particular, $\llbracket G\rrbracket$ is not boundedly acyclic when $G$ is not.